### v3

# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

Abstract.
LLM chains enable complex tasks by decomposing work into a sequence of subtasks. Similarly, the more established techniques of crowdsourcing workflows decompose complex tasks into smaller tasks for human crowdworkers. Chains address LLM errors analogously to the way crowdsourcing workflows address human error. To characterize opportunities for LLM chaining, we survey 107 papers across the crowdsourcing and chaining literature to construct a design space for chain development. The design space covers a designer’s objectives and the tactics used to build workflows. We then surface strategies that mediate how workflows use tactics to achieve objectives. To explore how techniques from crowdsourcing may apply to chaining, we adapt crowdsourcing workflows to implement LLM chains across three case studies: creating a taxonomy, shortening text, and writing a short story. From the design space and our case studies, we identify takeaways for effective chain design and raise implications for future research and development.
Three components: input task, design space, and resulting workflow. Input task is shortening text to specified lengths while preserving meaning. Design space flows through 3 main steps 1) Objectives which are to have a creative response while maintaining accurate actuality, 2) strategies, which are for response diversity, user guidance, and validation, and 3) tactics of: parallel generation, step to select focus, transform outputs, and evaluate candidates. After design space comes the resulting workflow. This workflow inputs a long text, finds areas to edit, fixes those areas, verifies those fixes, and outputs a shorter text with a length slider. The fix and verify steps have parallel generation.
1. Introduction
People use Large Language Models (LLMs) for assistance with both single-step tasks, like suggesting synonyms, and more complex multi-step tasks, like writing stories, editing text, and organizing ideas. Despite their widespread adoption, LLMs suffer from quality deficits like hallucinations (Dziri et al., 2022; Nair et al., 2023b; Maynez et al., 2020), brittleness to prompt changes (Gero et al., 2023; Arora et al., 2022; Zamfirescu-Pereira et al., 2023; Nie et al., 2022; Holtzman et al., 2021), and user interventions limited to prompting (Wu et al., 2023c; Parameswaran et al., 2023). These deficits are exacerbated for complex tasks, which require multiple steps in which errors can arise and propagate (Dziri et al., 2023). Unfortunately, many common use cases require multiple reasoning steps. For example, the use case of shortening text requires first deciding which portions of the text to edit and then determining how to revise each one (Bernstein et al., 2010).
To tackle complex tasks, recent research has turned to LLM chaining techniques. Chaining decomposes a task into multiple calls to an LLM, in which the output of one call affects the input to the next call (Wu et al., 2021). For example, when shortening text, an LLM chain could identify verbose sentences, edit each one, and propose outputs of variable lengths by composing multiple edits. Chaining has many benefits: this strategy can enable new abilities like long-form story writing (Yang et al., 2022; Mirowski et al., 2023), increase accuracy on logical reasoning questions (Xie et al., 2023; Zhang et al., 2022; Creswell and Shanahan, 2022), and enable greater transparency and debuggability (Huang et al., 2023b; Wu et al., 2022a; Reppert et al., 2023). Although initial evidence suggests that chaining can help mitigate LLM quality deficits, current solutions are imperfect (Nair et al., 2023a; Mirowski et al., 2023; Creswell and Shanahan, 2022; Park et al., 2022; Long et al., 2023; Josifoski et al., 2023). LLM chaining is nascent and challenging, with a seemingly intractable design space.
The challenge of decomposing complex work, however, is not new. In the more established field of crowdsourcing, researchers have faced a similar problem setup. Groups of people (requesters) outsource tasks to other people (crowdworkers) on online platforms. Crowdworkers have their own limitations: approximately 30% of crowdworkers’ outputs are low quality (Bernstein et al., 2010), small changes in instructions affect crowdworker responses (Kittur et al., 2012; McDonnell et al., 2016), and requesters have limited modes of interaction with crowdworkers (Salehi et al., 2017). To compensate for these limitations, crowdsourcing researchers have developed a variety of structured workflows. Crowdsourcing workflows decompose a task into a series of smaller microtasks for crowdworkers to complete. These workflows account for crowdworker limitations by reducing task complexity and leveraging redundancy, aggregation, and explicit validation (Bernstein et al., 2010; Cheng and Bernstein, 2015; Ambati et al., 2012; Zaidan and Callison-Burch, 2011).
Given the parallels between these fields, we investigate how crowdsourcing techniques may be adapted to design LLM chains (Figure 1). Selecting relevant strategies from over a decade of crowdsourcing research is challenging, especially as LLMs and crowdworkers have salient differences (Wu et al., 2023c). For example, crowdworkers may be slow to generate large text amounts, whereas LLMs may lack common sense and factual grounding, making them more sensitive to instructions and likely to “hallucinate” information.
To guide LLM chain development, we first construct a design space based on a systematic review of the crowdsourcing workflow and LLM chaining literatures. We analyze 107 papers and perform open coding to identify core design space dimensions. Our design space is organized into the objectives for which a designer can optimize a workflow (Figure 2) and the tactics used to build the workflow (Figure 3).
From this review, we also surface the strategies used to marshal and combine tactics to achieve objectives. For example, the objective to encourage creativity in outputs can be achieved by strategies that produce diverse responses, allow user guidance, and utilize adaptable architectures. We gather from the literature how tactics have been used in prior work to support strategies. For example, workflows and chains encourage diverse responses with redundant or communicative tactical architectures. Strategies constrain the space and evaluation of tactic choices for both crowds and LLMs (Figure 4).
To better assess the impact of similarities and differences between crowdworkers and LLMs on the implementation of strategies, we implement three case studies adapting crowdsourcing workflows to use LLMs. We chose workflows that enable direct manipulation of the output, an instance of the user guidance strategy for quality control that has been explored in crowdsourcing, but is underexplored in chaining. The case studies implement tasks that vary in their degrees of creativity, adapting Cascade (Chilton et al., 2013) to control the precision of a generated taxonomy, Soylent (Bernstein et al., 2010) to shorten text, and Mechanical Novel (Kim et al., 2017) to control the narrative elements of a fictional story.
We evaluate these chains in terms of the quality and controllability of their outputs. For example, compared to zero-shot prompting, chain-generated stories are preferred by crowdsourced raters and the story elements can be more precisely controlled. These case study implementations provide insights into how differences between LLMs and crowdworkers impact chain design to achieve the same objectives. In Mechanical Novel, for instance, we find that the LLM chain requires more explicit support than the crowdsourcing workflow to create creative and varied outputs.
Informed by both our design space and the three case studies, we discuss how differences between LLMs and crowdworkers impact effective tactic designs for LLM chains. For example, because LLM stochasticity differs from human diversity, tactic choices for LLMs must induce response diversity more explicitly and account for limitations in LLM self-validation. We find that our strategies help focus workflow adaptation and suggest areas of promising future work. Such areas include how the relative speed and reduced cost of LLMs relative to crowdworkers presents an opportunity to create higher quality chains, and we recommend improved tools and methods to augment this design process.
2. Related work
The bulk of our design space contribution is a literature review, so most related work has been discussed inline throughout the paper. Concurrent to this paper, a few other groups investigate the intersection of crowdsourcing and LLMs. One paper explores the impact of incorporating LLMs into crowdsourcing workflows on requester and crowdworker stakeholders (ALLEN et al., 2023). A second considers hybrid Crowd-LLM data annotation (Li, 2024). Another implements three strategies from crowdsourcing using LLMs for data processing tasks (Parameswaran et al., 2023). Closest to our work, the final paper has students implement six crowdsourcing workflows. The paper discusses the effectiveness of these workflows and the impact of the difference between LLM and crowdworker abilities on recommendations for chaining (Wu et al., 2023c). We build on this approach of providing recommendations based on case study implementations. We widen the scope of our recommendations by first integrating a broad survey of LLM chaining and crowdsourcing workflow literature into a design space and then using the case studies to explore open questions in the design space. This broader investigation enables us to build a design space and informs more comprehensive recommendations for navigating design decisions and for orienting future work. These recommendations span from higher-level guidance, such as adhering to the tradeoff between outcome quality and resource constraints, to lower-level suggestions, such as pursuing architecture-based methods for quality validation.
3. Design space methodology
We survey papers that involve the design and implementation of crowdsourcing workflows or LLM chains. Although crowdsourcing workflows and LLM chains have many overlapping concepts, the terms used to denote these concepts differ. In this paper, we refer to both LLMs and crowdworkers as “workers.” We use the term “workflow” to refer to both crowdsourcing workflows and LLM chains. We refer to intermediate steps of workflows or chains as “subtasks.” We use the term “user” to describe the person receiving workflow outputs (e.g., the “requester” in crowdsourcing terms).
3.1. Paper collection
We run two paper collection processes, one for the crowdsourcing literature and another for the LLM chaining literature. We use snowball sampling from the relevant citations in and of these papers, including every citation of the initial AI Chains paper (Wu et al., 2022b). We consider papers available as of August 2023.
Workflow literature. We collect papers through search of the keywords: crowd(sourcing) workflow, crowd(sourcing) pipeline, crowdsourcing complex work and microtask. We search the ACM CHI Conference on Human Factors in Computing Systems, ACM Conference on Computer-supported Cooperative Work and Social Computing, the ACM Symposium on User Interface Software and Technology, and the AAAI Conference on Human Computation and Crowdsourcing, as well as Semantic Scholar (Fricke, 2018) and Google Scholar.
Chaining literature. We collect papers through search of the keywords: LLM chain(ing), prompt chain(ing), LLM pipeline, and AI chain(s/ing). We search the International Conference on Learning Representations, the Conference on Neural Information Processing Systems, the International Conference on Machine Learning, the Association for Computational Linguistics proceedings, the Association for Computing Machinery proceedings, as well as Semantic Scholar (Fricke, 2018) and Google Scholar. We include preprints in our literature search.
Paper inclusion criteria. We include papers that implement, or support the implementation of, a workflow with at least two steps; i.e., workflows in which the outputs from one worker inform the input to another worker. We include papers that focus on building an interface or domain specific language for other designers to build workflows. We restrict our search to academic papers, rather than software like LangChain (LangChain, 2022). We exclude works that call multiple workers in parallel but never in a chain of steps (e.g., majority voting of labels). We also exclude techniques such as Chain of Thought (Wei et al., 2022) which issue multiple tasks to a worker in one step instead of incorporating multiple worker calls (Cheng et al., 2015). We include instances in which workers synchronously work on the same interface, viewing and editing each other’s outputs. Finally, the task at hand must have a user working towards some end goal, so we do not consider open-ended collaborative projects like Wikipedia. We make one exception for a paper that explicitly draws from crowdsourcing strategies to complete data aggregation tasks with LLMs (Parameswaran et al., 2023). Although they do not implement chains, the connection to crowdsourcing resonates with our design space goals.
Results. Our initial keyword search, conducted in August 2023, found 313 papers for workflows and 180 papers for chaining. One author determined whether papers were out of scope given our collection criteria, and any ambiguous cases were discussed among multiple authors. Our final set of papers resulted in 107 papers: 68 for workflows, and 39 for chains.
3.2. Coding strategy
We conduct a thematic analysis (Braun and Clarke, 2006), following the methodologies of similar design spaces in HCI (Lai et al., 2023; Bae et al., 2022; Shi et al., 2021). First, one author extracted 1) the outcomes (objectives) and 2) the elements (tactics) of the workflows from the paper set. We then employed iterative open coding on the extracted data, in which two authors separately coded each data extraction, sorted resulting codes into sub-categories, then discussed these categorizations. In three further rounds of iteration, the same authors refined these categorizations. All authors reviewed the resulting design space.
4. Design Space
Our design space has two parts: the objectives span the space of a designer’s goals, and tactics span the space of workflow implementations. These objectives are 1) promoting outcome qualities while respecting resource constraints, 2) increasing the accuracy and/or creativity of outputs, and defining the generality or specificity of the set of tasks that the designer wants the workflow to support. For the workflow tactics, workers complete various subtasks that are linked together into the workflow’s architecture.
Title of ”Objectives, what the designer wants the workflow to accomplish”. Three rows of text each with two objectives, examples, and a description. Row 1. A tradeoff between outcome quality (ex. transparency, no hallucination) and resource constraints (ex. effort, latency, expense). Description: ”Designers must work within resource constraints to improve outcome quality”. Row 2. A spectrum between accuracy (ex. math, question-answering) and creativity (ex. writing, prototyping). Description: Different tasks may require accuracy, creativity, or some combination thereof. Row 3. A spectrum between task specific (ex. only applied to one problem (e.g., translation workflow)) and task general (ex. can be applied to many problems(e.g., map-reduce)). Description: Workflows vary from supporting a specific task to supporting a general set of tasks.
4.1. Objectives in the design space
The objectives of a workflow design define the designer’s goals within several dimensions. Designers will design for some point along a outcome quality vs resource constraints tradeoff, in a range from accuracy to creativity, and in a range spanning between task specificity and generality. (Figure 2).
4.1.1. Outcome quality vs resource constraints
When working towards outcome qualities, designers must be mindful of resource constraints. Workflows can support quality outputs such as by improving accuracy and/or creativity (Kim et al., 2023b; Creswell and Shanahan, 2022; Lasecki et al., 2013; Reppert et al., 2023; Wu et al., 2021), transparency (Creswell et al., 2022; Creswell and Shanahan, 2022; Wu et al., 2023c; Reppert et al., 2023; Wu et al., 2023a; Kim et al., 2023b), debuggability (Wu et al., 2022a, 2021), and modularity (Huang et al., 2023b; LaToza et al., 2014; Lee et al., 2017; Retelny et al., 2014; Zharikova et al., 2023; Yao et al., 2023; Wu et al., 2023a).
Workflows can also address, although imperfectly, quality concerns. Some quality concerns are induced by the workers, such as hallucinations (Nair et al., 2023a; Mirowski et al., 2023; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023) and generative loops (Mirowski et al., 2023) in the case of LLMs, and fatigue (Dai et al., 2013; Alshaibani et al., 2020) in the case of human workers. Others are induced by the workflow itself, such as cascading errors (Wu et al., 2022a, 2021; Xie et al., 2023) and lack of coherency (Bernstein et al., 2010; Yang et al., 2022; Kim et al., 2017; Hahn et al., 2016; Mirowski et al., 2023; Wu et al., 2022a; Hahn et al., 2016). More concerns are induced by the broader context, such as the need to scale (Arora et al., 2022; Kulkarni et al., 2014, 2012), the workflow’s placement within a larger system (Pavel et al., 2014; Christoforaki and Ipeirotis, 2014; Cheng and Bernstein, 2015), unethical use of the workflow (Park et al., 2022), and the impact on human labor (Mirowski et al., 2023; Wu et al., 2023a), sense of ownership (Mirowski et al., 2023) and sense of agency (Kim et al., 2023b; Mirowski et al., 2023). However, mechanisms for improving output quality can be more expensive (Yao et al., 2023; Huffaker et al., 2020), take longer (Chang et al., 2017; Gouravajhala et al., 2018), and require more effort to create (Kittur et al., 2011; LaToza et al., 2014).
Expense. The expense of a workflow refers to paying crowdworkers or paying for an API call to an LLM. Workflows are often more expensive than simpler approaches (e.g., Chain of Thought prompting) (Zharikova et al., 2023; Yao et al., 2023; Wang et al., 2022; Du et al., 2023; Gero et al., 2023; Hahn et al., 2016; Huffaker et al., 2020; Lin et al., 2014a; Dai et al., 2010; Marcus et al., 2011a, b; Nair et al., 2023a). However, the cost is often less than hiring professionals (Ambati et al., 2012; Christoforaki and Ipeirotis, 2014; Zaidan and Callison-Burch, 2011; Liem et al., 2011; Fleming et al., 2023), pretraining models (Kim et al., 2023b; Levine et al., 2022; Yao et al., 2023; Huang et al., 2023b; Cobbe et al., 2021), or completing the task yourself (Zhang et al., 2022; Wang et al., 2023).
Latency. Another constraint is the time for the workflow to complete (Lasecki et al., 2015; Kulkarni et al., 2014; Marcus et al., 2011a, b). Many workflows have lower latency than a single person performing the same task, as work can be parallelized across multiple workers (Teevan et al., 2016; Lasecki et al., 2015; Chilton et al., 2014; Guo et al., 2016; Lee et al., 2017; André et al., 2013; Pietrowicz et al., 2013; Lin et al., 2014a; Gouravajhala et al., 2018; Kulkarni et al., 2014).
Effort. Workflows can help users complete tasks with less effort (Mirowski et al., 2023; Wu et al., 2023a; Kulkarni et al., 2014; Long et al., 2023). However, workflows can require high effort to create (Kulkarni et al., 2012; Hahn et al., 2016; Kittur et al., 2012) and oversee (Wu et al., 2022a; LaToza et al., 2014).
4.1.2. Accuracy vs creativity
One element of output quality of particular importance is if the workflow should support the accuracy or creativity of outputs. Many tasks’ definitions of output quality have elements of both accuracy and creativity.
Accuracy. Accuracy-focused tasks have one correct answer. Some crowdsourcing papers in this space include TurKit (Little et al., 2010b) and Second Opinion (Mohanty et al., 2019). Examples of these tasks include math problems (Zhou et al., 2022b; Liang et al., 2023; Du et al., 2023; Wu et al., 2023b; Madaan et al., 2023b), blurry text recognition (Little et al., 2010b, a), and matching images with the same content (Mohanty et al., 2019; Kobayashi et al., 2018). For these tasks, workflows can improve performance on benchmarks (Schick et al., 2022; Guo et al., 2019; Zhang et al., 2022; Cobbe et al., 2021; Li et al., 2023; Wu et al., 2023b; Eisenstein et al., 2022; McDonnell et al., 2016; Kobayashi et al., 2018), enable new tasks (Guo et al., 2017; Huffaker et al., 2020; Gouravajhala et al., 2018; Little et al., 2010b; Huang et al., 2023c), classify examples for which machine learning approaches struggle (Mohanty et al., 2019; Bernstein et al., 2010; Cheng and Bernstein, 2015; Alshaibani et al., 2020), improve instructional clarity (Bragg et al., 2018), and reduce hallucinations (Nair et al., 2023a; Mirowski et al., 2023; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023).
Creativity. Creative tasks have the widest degree of output freedom. Representative crowdsourcing papers for creative objectives include Flash Teams (Retelny et al., 2014), VisiBlends (Chilton et al., 2019), and Mechanical Novel (Kim et al., 2017). These papers tackle creative writing (André et al., 2014; Kim et al., 2017; Bernstein et al., 2010; Yao et al., 2023; Yang et al., 2022; Mirowski et al., 2023; Zhang et al., 2023a), idea generation (Little et al., 2010a; Kim et al., 2023b; Wu et al., 2023c), and prototyping (Lee et al., 2017; Kulkarni et al., 2014; Valentine et al., 2017; Retelny et al., 2014; Park et al., 2022). Prior work finds that on creative tasks, workflows can help achieve higher quality (André et al., 2014; Kim et al., 2017; Yao et al., 2023; Yang et al., 2022; Mirowski et al., 2023; Kim et al., 2023b; Lee et al., 2017; Park et al., 2022) and less constrained outputs (Liu et al., 2018; Hahn et al., 2016; Lee et al., 2017; Drapeau et al., 2016; Wang et al., 2023; Park et al., 2022). Workflows may enable novice crowdworkers (Pietrowicz et al., 2013) or users (Wang et al., 2023; Lin et al., 2014a; André et al., 2013; Wu et al., 2023c; Valentine et al., 2017; Chilton et al., 2019; Kim and Monroy-Hernandez, 2016) to perform a creative task with an otherwise high barrier of entry. Important qualities for creative tasks include coherence (Yao et al., 2023; Yang et al., 2022; Kim et al., 2017) and aligning to the user’s creative vision (Salehi et al., 2017). On creative tasks, a workflow is often better suited for creating and lengthening the output, rather than proofreading (Kim et al., 2017; Huang et al., 2017; Little et al., 2010b; Mirowski et al., 2023), so these outputs are sometimes seen as a starting point for editing, rather than a finished product (Kim and Monroy-Hernandez, 2016; Park et al., 2022; Long et al., 2023; Mirowski et al., 2023; Nebeling et al., 2016).
Both accuracy and creativity. Many tasks require some degree of accuracy and some degree of creativity. In some tasks, there is not one correct answer, but there are restrictions on what outputs are correct. Crowdsourcing papers for these tasks include Cascade (Chilton et al., 2013), Flock (Cheng and Bernstein, 2015), Revolt (Chang et al., 2017), and Storia (Kim and Monroy-Hernandez, 2016). Example tasks include categorizing data (categories must be accurate, but there are multiple options and unclear edge cases) (Bragg et al., 2013; Chilton et al., 2013; André et al., 2013; Chilton et al., 2014; Cheng and Bernstein, 2015; Chang et al., 2017), summarizing text (summaries must stay faithful, but have degrees of freedom in how to write the summary) (Kim and Monroy-Hernandez, 2016; Pavel et al., 2014; Bernstein et al., 2010; Nair et al., 2023a, b), and coding (programs must work, but there are multiple approaches to problem solving) (LaToza et al., 2014; Kulkarni et al., 2014; Wu et al., 2023a; Huang et al., 2023b, a). The goals of workflows for these tasks are to align with expert outputs (Kim et al., 2014; Liu et al., 2018; Chilton et al., 2013; Salehi et al., 2017; Park et al., 2022; Kim et al., 2023b; Fleming et al., 2023) or to enable users to choose among multiple correct possible output options. Using workflows for these tasks can result in an output that is human-understandable (Cheng and Bernstein, 2015), that better represents what would be important for people to know (Salisbury et al., 2017), and that surfaces the uncertainty among the different possible correct options (Chang et al., 2017).
4.1.3. Task specificity vs generality
Some workflows are highly specialized to the specifications of the task (e.g., visual blends (Wang et al., 2023; Chilton et al., 2019), masking faces in images (Alshaibani et al., 2020; Kaur et al., 2017), and programming (Huang et al., 2023b; LaToza et al., 2014)). Others generalize to a variety of tasks (Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Kulkarni et al., 2012; Wu et al., 2023a). For example, AutoGen (Wu et al., 2023a) uses the same conversation-based workflow for coding as well as chess games, math problems, and retrieval-based question answering. Task-specific workflows support fewer outcome types but can incorporate more tailored support than workflows that generalize to many tasks.
Title: Tactics, the building blocks of workflows. One row each for workers, subtasks, and architectures, with the category and (a short description). The workers row contains: Crowdworkers (people on an online platform), LLMs (generative large language models), user (the person deploying the workflow). The subtasks row contains: generate (create content), evaluate (judge content), transform (update content), focus (select content) and partition (split tasks). The architecture row contains: sequential (call LLMs in sequence), branching (diverge on paths), redundant (multiple workers repeat subtask), dynamic (Initialize parts of the workflow at runtime), and communicative (facilitate interaction among workers) Each subtask category is connected with a generic example instruction. Generate: ”Write a caption”. Evaluate: ”Is this answer grammatical?”. Transform: ”Improve the given idea”. Focus: ”Select regions for future work:. Partition: ”Divide the task into subtasks”.
4.2. Tactics in the design space
Designers use workflow tactics to implement workflows. Tactics need to consider the workers, the subtasks the workers complete, and the connections of subtasks into workflow architectures (Figure 3).
4.2.1. Workers
Workers may be crowdworkers, LLMs, or the user.
Crowdworkers. These workers are usually recruited from crowdworking platforms like Amazon Mechanical Turk or UpWork (formerly known as oDesk). Platforms allow for requests of worker qualifications, although these credentials are sometimes hard to validate (Ahmad et al., 2011; Zaidan and Callison-Burch, 2011; Valentine et al., 2017; Kulkarni et al., 2014). Methods like retainer models can reduce the latency of accessing crowdworkers on these platforms (Bernstein et al., 2011). Workers can also be recruited from the public, including community members (Mahyar et al., 2018), conference committees (Chilton et al., 2014; André et al., 2013), and online deaf communities (Bragg et al., 2021).
Crowdworkers have limitations that overlap with LLM limitations. Prior work suggests that approximately 30% of outcomes from Amazon Mechanical Turk are low quality (Bernstein et al., 2010), and may include spam or off-topic responses (Butler et al., 2017; Lasecki et al., 2013; Alshaibani et al., 2020; Huang et al., 2018). Crowdworkers also struggle with following long and complex instructions, as people can overly focus on the beginning or end of a given text (Hahn et al., 2016; Chilton et al., 2013; Lasecki et al., 2015; Chen et al., 2019). Crowdworkers can be biased; for example, in one study, workers performed significantly better in identifying photos of their own race compared to other races (Mohanty et al., 2019).
Crowdworkers have unique limitations. Anchoring bias can lead to less creativity in generation or divergent thinking (Wu et al., 2022b, 2023c). Since workers are paid when they complete a task, there is incentive to work quickly even if that means satisficing (Bernstein et al., 2010; Parameswaran et al., 2023; Hata et al., 2017). Additionally, since rejected work can hurt access to future jobs on a platform (Kim et al., 2017), workers may overcompensate to a detrimental level in an attempt to make their efforts clear (Kim et al., 2014; Bernstein et al., 2010). Crowdworkers may fatigue (Alshaibani et al., 2020; Dai et al., 2013), and they may lack task-relevant knowledge and delete jargon (Nebeling et al., 2016; Bernstein et al., 2010). On synchronous tasks, other limitations become prevalent. Social blocking can make workers so cautious to change others’ work that not enough work gets done (Lasecki et al., 2015; Hahn et al., 2016), and territoriality can motivate edit wars in which workers’ attempts to establish their own work over others degrade the final product quality (Valentine et al., 2017; André et al., 2014). Crowdworkers also may be affected by diffusion of responsibility and social loafing in which there is incentive to wait for others to do tasks (Lasecki et al., 2015; André et al., 2014).
Crowdworkers have their comparative strengths as well, including the ability to perform high-level structural evaluation and to quickly identify outliers (Huang et al., 2015; Luther et al., 2015). People are also currently better at selectively choosing among subtasks than LLMs (Wu et al., 2023c).
LLMs. There are many possible LLMs of many sizes and expenses (Bursztyn et al., 2022). Models have differing strengths and respond differently to prompts (Mirowski et al., 2023; Wu et al., 2023c; Parameswaran et al., 2023; Liang et al., 2023). LLM outputs can be adjusted through parameter changes (e.g., temperature) (Park et al., 2022; Parameswaran et al., 2023), fine-tuning (Creswell et al., 2022; Creswell and Shanahan, 2022; Cobbe et al., 2021; Li et al., 2023; Bursztyn et al., 2022), and expressing roles through prompts (Park et al., 2022; Du et al., 2023).
Some LLM limitations overlap with those of crowdworkers. LLMs can provide erroneous outputs: they are known to confidently output incorrect answers (Parameswaran et al., 2023; Wang et al., 2023; Du et al., 2023) or hallucinations (Nair et al., 2023a; Mirowski et al., 2023; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023) and for struggling with self-consistency and logical reasoning (Creswell et al., 2022; Creswell and Shanahan, 2022; Yao et al., 2023; Parameswaran et al., 2023; Liang et al., 2023; Xie et al., 2023). LLMs can ignore large fractions of long prompts or focus mostly on recent parts of the given context (Parameswaran et al., 2023; Du et al., 2023; Wu et al., 2023c; Madaan et al., 2023b). LLMs also output biased results, a widespread concern (Park et al., 2022; Wu et al., 2023a; Huang et al., 2023a; Long et al., 2023). For example, work on screenplay generation found stereotypical outputs, with multiple instances of gender biases (Mirowski et al., 2023). Furthermore, LLM training on human data increases the risk of accidental plagiarism (Mirowski et al., 2023).
LLMs also have unique limitations. On creative tasks, LLMs sometimes produce outputs that are perceived as robotic (Long et al., 2023) and that can lack the understanding, nuance, and subtext that people can provide (Mirowski et al., 2023). Furthermore, models can get stuck in generative loops, degenerating text quality (Parameswaran et al., 2023). LLMs also hallucinate more than crowds (Nair et al., 2023a; Mirowski et al., 2023; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023) and may struggle to change their outputs through self-reflection (Liang et al., 2023).
LLMs have comparative strengths as well. They may be better at divergent thinking (Wu et al., 2023a, c; Liang et al., 2023) and comparison-based instructions (Wu et al., 2022b, 2023c) than crowdworkers, who are affected more by anchoring bias. LLMs may also be better than crowdworkers at tasks requiring recollection of large amounts of data (Parameswaran et al., 2023).
User. Finally, users of the workflow can also be involved to revise or guide the process. Users can be novices using the workflow to adapt to a steep learning curve (Kulkarni et al., 2014; Pietrowicz et al., 2013; Lin et al., 2014a; Chilton et al., 2019; Wang et al., 2023), or experts using the workflow as a support tool (Mirowski et al., 2023; Nebeling et al., 2016; Mohanty et al., 2019). Involving users in completing subtasks grants them greater control at the expense of higher effort (Stretcu et al., 2023). However, even with user involvement, prior work finds mixed results on how well tools retain the user’s sense of agency and ownership (Kim et al., 2023b; Mirowski et al., 2023).
4.2.2. Subtasks
Subtasks are singular tasks in a workflow. Although highly varied, they follow several thematic functions: to generate, evaluate, transform, focus, and partition.
Generate. Most workflows involve at least one step in which workers generate content. This generation may be writing words or code (Pavel et al., 2014; Weir et al., 2015; Zhang et al., 2022; Li et al., 2023; Zhang et al., 2023a) brainstorming and ideation (Chilton et al., 2019; Lasecki et al., 2013; Mahyar et al., 2018; Huang et al., 2018; Long et al., 2023; Yao et al., 2023), or answering a question or solving a problem (Kobayashi et al., 2018; Wu et al., 2023b; Xie et al., 2023; Eisenstein et al., 2022). Generation tasks may also ask for labels or annotations on given content (Guo et al., 2017, 2019; Chilton et al., 2014; Chang et al., 2017; Bragg et al., 2013), to translate between languages (Ambati et al., 2012; Zaidan and Callison-Burch, 2011), to contribute to a conversation with other workers (Mahyar et al., 2018; Wang et al., 2022; Liang et al., 2023; Du et al., 2023), to demonstrate a behavior (Bragg et al., 2021; Lee et al., 2017), or to source content from external sources (Kim et al., 2014; Chilton et al., 2019; Hahn et al., 2016; Kim and Monroy-Hernandez, 2016; Pietrowicz et al., 2013; Huang et al., 2015; Luther et al., 2015; Christoforaki and Ipeirotis, 2014). Some subtasks also require an explanation along with the generated content (Chang et al., 2017; Kim et al., 2017; Cheng and Bernstein, 2015; Drapeau et al., 2016; Schick et al., 2022; Gero et al., 2023; Fleming et al., 2023).
Evaluate. One of the most common evaluation tasks is choosing the best among multiple possible options (Guo et al., 2017; Weir et al., 2015; Guo et al., 2019; Cheng and Bernstein, 2015; Kittur et al., 2012; Liu et al., 2018). Other subtasks may include rating and ranking among multiple options (Pavel et al., 2014; Salehi et al., 2017; Marcus et al., 2011a; Huffaker et al., 2020; Willett et al., 2012; Mahyar et al., 2018) or a pairwise comparison (Mohanty et al., 2019; Little et al., 2010b, a; Marcus et al., 2011a, b; Parameswaran et al., 2023; Bursztyn et al., 2022). Evaluation subtasks may also determine if another round of generation needs to occur (Wu et al., 2022b; LaToza et al., 2014; Wang et al., 2022; Liang et al., 2023).
Workflows can decide what moves forward through majority vote (Huang et al., 2017; Dai et al., 2010; Lasecki et al., 2015; Liu et al., 2018; Kim et al., 2014; Guo et al., 2019), other aggregations of voting tasks (e.g., expectation maximization (Drapeau et al., 2016; Parameswaran et al., 2023), weighted mean (Mohanty et al., 2019), weak supervision (Arora et al., 2022)), returning the highest ranked or rated items (Pavel et al., 2014; Huffaker et al., 2020), or automated methods. These automated methods include measuring agreement from the generation step (Liu et al., 2018; Hahn et al., 2016; Huffaker et al., 2020; Chang et al., 2017; Huang et al., 2015; Nair et al., 2023b; Bernstein et al., 2011) and using non-AI checks like semantic embeddings and rule-based heuristics (Parameswaran et al., 2023; Wang et al., 2023; Ambati et al., 2012). There are also subtasks that seek critiques and reviews from the user (Nebeling et al., 2016; Schick et al., 2022) or another worker (Kim et al., 2017; Kulkarni et al., 2014; Valentine et al., 2017; Christoforaki and Ipeirotis, 2014; Mahyar et al., 2018; Schick et al., 2022; Parameswaran et al., 2023; Madaan et al., 2023b; Liang et al., 2023). These subtasks may ask for feedback or verify if errors or other qualities exist (Huang et al., 2023b; Bragg et al., 2013; Christoforaki and Ipeirotis, 2014; Chilton et al., 2019; Weir et al., 2015; Huang et al., 2015; Kim et al., 2014; Gero et al., 2023; Huang et al., 2017; Bernstein et al., 2010; Yang et al., 2022; Wu et al., 2023c; Kittur et al., 2011; Liang et al., 2023; Kulkarni et al., 2012; André et al., 2013; Butler et al., 2017).
Transform. Other subtasks transform given data by merging or improving it. Some subtasks merge multiple streams of incoming content. This merge could be algorithmic (Alshaibani et al., 2020; Ahmad et al., 2011; Alshaibani and Quinn, 2021; Teevan et al., 2016; Chilton et al., 2019; Kittur et al., 2011) or accomplished by the worker (Kittur et al., 2011; Wu et al., 2023c, 2022b; Lee et al., 2017; LaToza et al., 2014; Ambati et al., 2012; Teevan et al., 2016; Kittur et al., 2012; Wu et al., 2023c; Pietrowicz et al., 2013; Luther et al., 2015; Kulkarni et al., 2012). Other subtasks improve content generated by previous steps, sometimes by incorporating feedback from evaluation steps. Improving on previously generated content can come in many forms: updating the format of the content (Huang et al., 2023b; Wu et al., 2022b; Eisenstein et al., 2022; Wang et al., 2022; Arora et al., 2022), pruning bad results (Liu et al., 2018; Gero et al., 2023), adding missing results (Huang et al., 2023b; Liu et al., 2018; Weir et al., 2015), correcting errors (Huang et al., 2023b; Yang et al., 2022; Drapeau et al., 2016; Kobayashi et al., 2018; Gouravajhala et al., 2018), and performing direct improvements (Madaan et al., 2023b; Hahn et al., 2016; Mahyar et al., 2018; Lee et al., 2017; Huffaker et al., 2020; Nebeling et al., 2016; Christoforaki and Ipeirotis, 2014; Little et al., 2010b; Liem et al., 2011; Dai et al., 2010; Schick et al., 2022; Kim et al., 2017; Bernstein et al., 2010; Guo et al., 2019; McDonnell et al., 2016).
Focus. One set of subtasks looks to select the attention of the upcoming tasks within a broader context. For example, this step may find patches of the input that need editing, extract information from context, or focus on common keywords (Bernstein et al., 2011; Huffaker et al., 2020; Gero et al., 2023; Creswell and Shanahan, 2022; Creswell et al., 2022; Nair et al., 2023b; Kim et al., 2017).
Partition. Some subtasks transform a task by breaking it up into smaller pieces (Kittur et al., 2011; LaToza et al., 2014; Kulkarni et al., 2012; Wu et al., 2023c; Zhou et al., 2022b; Valentine et al., 2017; Wu et al., 2022b; Huang et al., 2023c). For example, the Divide step in the Turkomatic workflow asks workers to divide a given subtask into simpler subtasks. The workflow calls this step recursively until the subtask complexity is valued (by a separate crowdworker) at a specified cost (Kulkarni et al., 2012).
Three columns headed by ”If…”, ”Then…” and ”With…”. The ”If” column contains different possible challenges relating to achieving objectives. The ”Then” column lists all of the strategies. There are arrows connecting the ”If” column to the ”Then” column, with arrows from ”navigating a large creative space” objectives challenge to the adaptive architectures, diverse responses, and user guidance strategies. The ”With” column lists all of the tactics. There are arrows connecting the ”Then” column to the ”With” column. There are many arrows, and they correspond to the tactics used for strategies as defined in the text and other figures.
4.2.3. Workflow architecture
Workflows combine subtasks into larger architectures that may be sequential, branching, redundant, dynamic, or communicative. One workflow may employ multiple architectural patterns.
Sequential architectures order subtasks in a chain in which the output of one worker passes forward to another worker (Wang et al., 2023; Nair et al., 2023a; Wu et al., 2022b; Huang et al., 2023a; Zhou et al., 2022b; Huang et al., 2023b; Reppert et al., 2023; Gero et al., 2023; Wu et al., 2023c; Long et al., 2023; Eisenstein et al., 2022).
Branching workflows apply conditionals to choose different execution paths based on the outputs of prior steps (Huang et al., 2018, 2023b; Kittur et al., 2012; Wu et al., 2022a; Zharikova et al., 2023).
Redundant architectures require different workers to complete the same task in parallel or iteratively. In parallel generation, multiple workers perform the same subtask separately (Weir et al., 2015; Guo et al., 2019; Cobbe et al., 2021; Butler et al., 2017; Lin et al., 2014a; Alshaibani and Quinn, 2021; Huang et al., 2017). This subtask is often, but not always, followed by an aggregation or evaluation subtask to merge or select among the different parallel generations. Iterative architectures run iteratively over a changing input. For example, iterative improvement takes the output of a subtask and iteratively gives it to a series of workers to improve, sometimes with interspersed decision tasks choosing between the current and prior iterations (Luther et al., 2015; André et al., 2014; LaToza et al., 2014; Goto et al., 2016; Dai et al., 2013; Lasecki et al., 2015; Willett et al., 2012). The number of iterations can be determined to be a fixed number of steps (Wu et al., 2023c; Kim et al., 2017; Huffaker et al., 2020; Little et al., 2010b; Zaidan and Callison-Burch, 2011), by including a subtask to check for sufficiency (Madaan et al., 2023b; Creswell et al., 2022; Creswell and Shanahan, 2022; Nair et al., 2023b; Wang et al., 2022; Wu et al., 2023b; Liang et al., 2023; Salehi et al., 2017; Kulkarni et al., 2014; Guo et al., 2016), until the user stops calling iterations (Cheng and Bernstein, 2015; Reppert et al., 2023; Kulkarni et al., 2014; Chilton et al., 2019; Willett et al., 2012), until the outputs converge (Schick et al., 2022; Liem et al., 2011; Du et al., 2023; Gero et al., 2023; Liem et al., 2011), until all quality checks are passed (Kim et al., 2017; Bragg et al., 2013; Chilton et al., 2013; Marcus et al., 2011a, b; Yang et al., 2022; Teevan et al., 2016), from increasing complexity levels (Alshaibani and Quinn, 2021; Alshaibani et al., 2020; Bursztyn et al., 2022), and through decision-theoretic approaches (Park et al., 2022; Dai et al., 2010).
Dynamic architectures support changing the workflow architecture as it is executed. Map-reduce is a dynamic architecture in which workers break down a task recursively until the subtasks are an appropriate size (Wu et al., 2023c; Liem et al., 2011; Kittur et al., 2011; Kim and Monroy-Hernandez, 2016; Kulkarni et al., 2012; Huang et al., 2023c; Kaur et al., 2017; Valentine et al., 2017; Liem et al., 2011). The literature defines such workflows as a hierarchy (Huang et al., 2023b; Kittur et al., 2011). Other dynamic workflow architectures include doing beam search across multiple options (Xie et al., 2023; Yao et al., 2023; Creswell and Shanahan, 2022) and allowing for workers to flexibly create tasks on a shared to-do list before choosing which of those to complete (Chilton et al., 2019; Nebeling et al., 2016; Mahyar et al., 2018; Zhang et al., 2012; Kittur et al., 2011; LaToza et al., 2014). Some other architecture methods may have dynamic elements, such as parallelizing an input into chunks and running the workflow for each chunk (André et al., 2013; Gouravajhala et al., 2018; Zaidan and Callison-Burch, 2011; Kim et al., 2014; Pavel et al., 2014; Kaur et al., 2017; Bernstein et al., 2010) or dynamically determining the number of iterative steps (Chen et al., 2019; Chilton et al., 2014; Madaan et al., 2023b; Zhou et al., 2022b; Nair et al., 2023b).
Communicative architectures can facilitate debate and conversation among workers (Du et al., 2023; Teevan et al., 2016; Wu et al., 2023a; Salisbury et al., 2017; Drapeau et al., 2016; Chen et al., 2019; Chang et al., 2017; Kobayashi et al., 2018; Liang et al., 2023) and allow workers to return feedback to the user (Kulkarni et al., 2012; Lasecki et al., 2015; Nebeling et al., 2016). Some workflows create roles for different workers, often based on roles held in real-life organizations (e.g., journalists, software development teams, etc.) (Valentine et al., 2017; Agapie et al., 2015; Retelny et al., 2014; Park et al., 2022). Some papers integrate these roles into a hierarchy of communication, with managers, midlevel managers, and different forms of communication between different levels of the hierarchy (Agapie et al., 2015; Retelny et al., 2014). Communication may be restrained through subtasks (Chang et al., 2017; Chen et al., 2019), free-form though email (Agapie et al., 2015; Valentine et al., 2017), or enabled through a shared interface (Lasecki et al., 2015; Lee et al., 2017; Nebeling et al., 2016; Kulkarni et al., 2014; Mahyar et al., 2018; Lasecki et al., 2013). Communication among workers may be synchronous or asynchronous. Synchronous workflows help enable collaboration (André et al., 2014; Chilton et al., 2014; Salisbury et al., 2017; Chilton et al., 2019) and can reduce latency (Bernstein et al., 2011). With asynchronous workflows or workflows with limited communication architectures, a working memory or state management can help workers access progress reports, what other workers think are good next steps, and the overarching goals (Valentine et al., 2017; Hahn et al., 2016; Lasecki et al., 2013; LaToza et al., 2014; Zhang et al., 2012; Hahn et al., 2016; Yang et al., 2022; Nair et al., 2023b; Wu et al., 2023a).
5. Strategies integrating the design space
Choosing tactic designs to support objectives is a non-trivial task, as combining different tactic types and implementations supports a huge range of possible workflows. However, through the thematic coding process we surfaced strategies by which tactics are used to support objectives (Figure 4). Specifically, these strategies (5.1) encourage diverse responses, (5.2) validate outputs from subtasks, (5.3) enable user guidance, (5.4) utilize adaptable architectures, (5.5) improve worker recruitment, (5.6) align a subtask’s required skills with a worker’s capabilities, (5.7) promote global context, and (5.8) ensure outputs meet a predetermined quality thresholds.
We outline the strategies that surfaced from the literature in a manner adapted from design patterns (Alexander, 1977; Gamma et al., 1994) used in architecture and software engineering. For each, we define the strategy, explain the effects on objectives that motivate using the strategy, and how to implement the strategy with tactics. We then discuss secondary benefits, tradeoffs, examples from the literature, and connections to other strategies. While we note some differences between the crowdsourcing and chaining literatures throughout this section, we leave the bulk of such discussion to Section 7.
5.1. Diverse responses
To improve accuracy or creativity, source multiple responses to the same subtask with parallel, iterative, or debate based workflows (Figure 5).
Motivation. Individual workers can give low quality outputs, affecting the quality of the overall output. Designers may want to improve robustness to low quality outputs, outliers, and spammers to support the accuracy of outcomes (Liem et al., 2011; Du et al., 2023; Zaidan and Callison-Burch, 2011; Huang et al., 2018; Lasecki et al., 2013; Huffaker et al., 2020; Reppert et al., 2023; Ambati et al., 2012; Hahn et al., 2016). For creative tasks, they may want to encourage divergent and diverse thinking (Liang et al., 2023; Wu et al., 2023a; Luther et al., 2015; Lasecki et al., 2015; Huang et al., 2017; Nebeling et al., 2016; Teevan et al., 2016).
Implementation. Getting responses from multiple workers on a subtask can help support these goals. Crowdsourcing workflows and LLM chaining source diverse responses differently. Crowdsourcing does so by querying different people. Chaining leverages the stochastic nature of LLMs (e.g., with a higher temperature parameter setting) or or uses different models or prompt variants (including role-based prompts) to source a variety of responses (Park et al., 2022; Wu et al., 2023c; Arora et al., 2022). In some cases one can automatically calculate the optimal mix of LLMs for a task (Madaan et al., 2023a).
Redundant architectures are the most direct way of sourcing diverse responses, but communicative architectures, such as those that structure debates, can source diverse responses as well. Some studies in crowdsourcing systematically study the impact of different architectures (e.g., parallel vs iterative) and architecture parameters (e.g., the number of iterations) on the diversity and quality of the outcomes (Goto et al., 2016; Little et al., 2010a; Kittur et al., 2011; André et al., 2014). However, it is unknown if these findings transfer to similar tactic choices with LLMs.
Secondary benefits. There are some benefits of diverse responses that are solely talked about in the crowdsourcing literature, as they are dependent on having different people. Crowdsourcing brings up the benefit of having different viewpoints (Chen et al., 2019; Ambati et al., 2012; Lasecki et al., 2013), filling in each other’s deficits to create a more rounded final product (Kim and Monroy-Hernandez, 2016; Zhang et al., 2012), and sourcing perspectives from expert crowdworkers (Chilton et al., 2014) or from workers with various skill levels (Mahyar et al., 2018).
Consequences. Sourcing diverse responses is inherently more costly because multiple workers complete the same subtask. Merging multiple responses can induce outcome quality errors: coherency for parallel architectures (Wu et al., 2022a) and cascading errors for iterative architectures (Goto et al., 2016).
Examples. Examples of the diverse responses strategy include: Apparition (Lasecki et al., 2015), chains for multi-agent debate (Liang et al., 2023), and this exploration of ”human computation processes” (Little et al., 2010a).
Related strategies. Agreement among diverse responses can serve as validation. Setting quality thresholds can impact the number of diverse responses requested.
Title. Diverse responses, to improve accuracy or creativity, source multiple responses to the same subtask with parallel, iterative, or debate based workflows. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Robust to low quality responses, Coherency and cascading errors), Constraints (increase expense and latency), Accuracy (Robust to inaccurate outputs), Creativity (Improve with divergent thinking). Section 2: tactics used. Workers (LLMs, crowdworkers), Subtasks (all), Architectures (redundant, communicative).
5.2. Validation
To improve quality of the overall output, validate and correct subtask outputs with deterministic, architectural, or worker-based checks (Figure 6).
Title. Validation, to improve quality of the overall output, validate and correct subtask outputs with deterministic, architectural, or worker-based checks. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Robust to low quality responses), Constraints (increase expense and latency), Accuracy (check for correctness), Creativity (confirm elements of quality exist), Task specific (can validate specific qualities). Section 2: tactics used. Workers (LLMs, crowdworkers, user), Subtasks (evaluate, transform), Architectures (sequential, redundant, communicative).
Motivation. Individual workers can give low quality responses. Unaddressed, these can lead to cascading errors through the workflow (Wu et al., 2022a; Goto et al., 2016; Little et al., 2010b; Bernstein et al., 2010; Little et al., 2010a). Therefore, higher quality subtask outcomes throughout the workflow can have myriad benefits on the overall outcome quality for both accuracy and creativity focused tasks, increasing robustness to low-quality worker responses (Ambati et al., 2012; Bernstein et al., 2010).
Implementation. The validation process is often sequential, in which content is first generated before being validated (Gero et al., 2023). In parallel and iterative redundant workflows, errors from prior passes can be found and fixed or rejected (Zaidan and Callison-Burch, 2011; Hahn et al., 2016) and differing opinions can be merged (Kittur et al., 2011; Luther et al., 2015) through transform and evaluate subtasks (Wu et al., 2023a, c; Zharikova et al., 2023; Kulkarni et al., 2012; Ambati et al., 2012; Schick et al., 2022). Users can break ties (Pietrowicz et al., 2013) and intervene if there are cascading errors (Kulkarni et al., 2012). Non-worker methods, such as taking the majority vote of the outputs of redundant evaluate subtasks can also be effective. Automated methods of flagging malicious workers, gold-standard validation steps, or requiring a certain agreement threshold among workers within a redundant architecture can also provide validation (Lasecki et al., 2015; Parameswaran et al., 2023; Huffaker et al., 2020; Liem et al., 2011).
Debate-based communicative workflows can promote reflection on disagreements (Kobayashi et al., 2018; Chen et al., 2019; Du et al., 2023; Liang et al., 2023). Using different models and roles in debate shows promise (Du et al., 2023), as an LLM judge in a debate was found to more likely agree with its same underlying model (Liang et al., 2023). Other workflows use adaptable architectures to make all production always editable so workers can fix errors in a adaptable manner (Liu et al., 2018; LaToza et al., 2014).
Consequences. Costs increase when validation involves more calls to workers. The literature shows mixed results for an LLM’s ability to validate its own responses in evaluate subtasks. Some studies suggest that adding a self-validation step improves outcomes (Madaan et al., 2023b; Xie et al., 2023; Gero et al., 2023; Yao et al., 2023), while others suggest that LLMs are likely to agree with their previous outputs, even when incorrect, and propose debate-based communicative workflows as an alternative (Liang et al., 2023; Du et al., 2023).
Examples. Examples of the validation strategy include: the “Verify” step of Soylent (Bernstein et al., 2010), a Dual-Pathways structure for transcription (Liem et al., 2011), and self-evaluation guided reasoning (Xie et al., 2023).
Related strategies. Validation often involves considering the agreement among diverse responses.
5.3. User guidance
To help the user explore the design space and align creative outcomes to their vision, incorporate their input into the workflow at runtime (Figure 7).
Title. User Guidance. To help the user explore the design space and align creative outcomes to their vision, incorporate their input into the workflow at runtime. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Users learn through interaction), Constraints (Increased effort), Accuracy (Review uncertainties), Creativity (Guide the creative vision), Task general (Can predefine less before runtime). Section 2: tactics used. Workers (user), Subtasks (evaluate, focus), Architectures (redundant, dynamic, communicative).
Motivation. Users may want to use the workflow to reduce the effort of exploring of the design space, to align the workflow’s output to their creative vision, or some combination thereof (Nebeling et al., 2016; Willett et al., 2012; Kulkarni et al., 2014; Park et al., 2022). If a process includes data that is inherently uncertain it could benefit from more explicit guidance by the user (Chang et al., 2017; Willett et al., 2012; Cheng and Bernstein, 2015).
Implementation. This strategy incorporates the user into a critical role within the workflow. Users can guide outcomes by dynamically answering questions and providing feedback (Salehi et al., 2017) or monitor dynamic workflows to catch cascading errors (Kulkarni et al., 2012) with evaluate or transform subtasks. Users can be called upon when necessary with a “panic button” that other workers can invoke (Kulkarni et al., 2012), or automatically when ties occur (Pietrowicz et al., 2013). With a shared communicative interface between users and all workers, users can intervene when needed and make real-time suggestions for prototyping directions (Nebeling et al., 2016; Lasecki et al., 2015). Users may review at intermediate steps to edit or regenerate sections of the workflow in evaluate or focus subtasks (Park et al., 2022; Kim et al., 2023b; Zhang et al., 2023a; Kulkarni et al., 2014). A redundant iterative architecture may also include multiple iterations with user revisions between each one (Willett et al., 2012; Cheng and Bernstein, 2015; Kulkarni et al., 2014; Park et al., 2022; Kim et al., 2023b). Sometimes disagreements raised with redundant architectures may not need to be resolved within the workflow and can instead support an explorable interface for users to directly manipulate (Willett et al., 2012; Chang et al., 2017; Cheng and Bernstein, 2015; Mohanty et al., 2019). They can sort and filter the returned data (Willett et al., 2012; Cheng and Bernstein, 2015; Mohanty et al., 2019), use a slider to control output characteristics like text length (Bernstein et al., 2010), and consider category labels at multiple levels of precision (Chang et al., 2017).
Crowdsourcing workflows integrate the user in more varied ways than LLM chaining, including providing results without aggregation that the user can sort and filter (Lin et al., 2014a; Mohanty et al., 2019), user guidance for prototyping (Cheng and Bernstein, 2015; Kulkarni et al., 2014; Retelny et al., 2014), and worker-initiated user revision (Salehi et al., 2017).
Secondary benefits. Incorporating the user into the process at runtime can reduce the amount of pre-definition of workflow steps needed, making workflows more generally applicable (Nebeling et al., 2016; Kulkarni et al., 2014).
Consequences. These benefits come at the cost of more effort for the user than an entirely automatic workflow.
Examples. Examples of the user guidance strategy include: Second Opinion (Mohanty et al., 2019), Social Simulacra (Park et al., 2022), and WearWrite (Nebeling et al., 2016).
Related strategies. Incorporating user guidance can aid with validation.
5.4. Adaptable architectures
To build workflows that require less effort to design and that generalize to many inputs, use adaptive architectures that build themselves at runtime (Figure 8).
Title. Adaptable Architectures. To build workflows that require less effort to design and that generalize to many inputs, use adaptive architectures that build themselves at runtime. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Cascading errors and infinite loops, Less transparency and consistency), Constraints (Reduce expense and latency), Creativity (Adapt to edge cases), Task general (Adaptable for task generality, Request tasks with no pre-definition, Decomposable input required). Section 2: tactics used. Subtasks (focus, partition), Architectures (branching, dynamic, communicative).
Motivation. It is challenging and requires high effort to design a rigid and pre-defined workflow that can effectively generalize to many tasks or that can accommodate for many edge-cases, especially in creative tasks (Kim and Monroy-Hernandez, 2016; Valentine et al., 2017; Kittur et al., 2011; Kulkarni et al., 2012). Instead, designers may want workflows that are adaptable and responsive (Valentine et al., 2017; Kim and Monroy-Hernandez, 2016; Kulkarni et al., 2014; Chen et al., 2019; Wu et al., 2023a; Nair et al., 2023b; Creswell and Shanahan, 2022) and to support a variety of inputs without the effort of explicit pre-definition (Kittur et al., 2011; Kulkarni et al., 2012; Creswell et al., 2022; Creswell and Shanahan, 2022).
Implementation. Adaptable workflows are not rigid in their progression of subtasks but instead are flexible and generate parts of the workflow architecture at runtime (Valentine et al., 2017; Kim and Monroy-Hernandez, 2016; Kulkarni et al., 2014; Chen et al., 2019; Kittur et al., 2011; Huang et al., 2023c, b). Workflows can incorporate subtasks defined by the crowd with partition and focus subtasks in dynamic and communicative architectures (Ahmad et al., 2011; Zhang et al., 2012; Agapie et al., 2015; Retelny et al., 2014; Reppert et al., 2023). Changes in these adaptable workflows may be made by the worker (LaToza et al., 2014; Valentine et al., 2017; Retelny et al., 2014), user (Retelny et al., 2014; Nebeling et al., 2016; Valentine et al., 2017; Kulkarni et al., 2012), or an algorithm (Bragg et al., 2013; Chen et al., 2019; Dai et al., 2013; Goto et al., 2016). Although branching architectures are predefined, they also allow for adaptability in which different inputs take different paths.
Consequences. However, there are tradeoffs. Adaptable workflows can lead to poor quality worker-created subtasks and instructions, causing cascading errors and infinite loops (Kulkarni et al., 2012). They also are restricted to tasks with an easily decomposable input (Kittur et al., 2011).
Chaining papers highly emphasize the importance of transparency (Creswell et al., 2022; Creswell and Shanahan, 2022; Reppert et al., 2023; Wu et al., 2022a; Reppert et al., 2023; Gero et al., 2023; Wu et al., 2021, 2023a), increasing the risk of using adaptable workflows which may cause unforeseeable errors as each step is not predefined. The degree to which adaptable workflows degrade this transparency is yet to be explicitly explored.
Some effects of adaptable workflows on worker experiences mattered for crowdsourcing, but do not matter for chaining. Crowdsourcing work has allowed workers to chose their own tasks as an incentive to improve task performance (Chilton et al., 2019; Nebeling et al., 2016; Mahyar et al., 2018; Zhang et al., 2012; Kittur et al., 2011; LaToza et al., 2014), but the need to incentivize workers does not apply to chaining. One crowdsourcing study that found users were frustrated with too much adaptability, an emotion that does not apply to LLMs (Valentine et al., 2017).
Title. Worker recruitment, To improve outcome quality at each step, curate a high quality worker pool by filtering workers by their skills and improving worker capabilities over time. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Higher quality subtask responses), Constraints (Increase cost and effort of filtering). Section 2: tactics used. Workers (LLMs, crowdworkers).
Examples. Examples of adaptable architectures include: map-reduce (Kittur et al., 2011), Flash Teams (Retelny et al., 2014), and microtask programming (LaToza et al., 2014).
Related strategies. Adaptable architectures must support the dynamic creation of aligned subtasks that have well defined context and worker instructions. Adaptable architectures can also dynamically change course once quality thresholds are met.
5.5. Worker recruitment
To improve outcome quality at each step, curate a high quality worker pool by filtering workers by their skills and improving worker capabilities over time (Figure 9).
Motivation. As workers power the workflow, workers that can provide higher-quality responses can help support outcome quality (Bragg et al., 2018; Liu et al., 2016).
Implementation. In crowdsourcing, recruiting high-quality workers can mean filtering crowdworkers by pre-specified qualifications or using platforms with more specialized workers such as Upwork (Valentine et al., 2017). For LLM chains, model choice and finetuning impact worker abilities (Cobbe et al., 2021; Bursztyn et al., 2022; Creswell et al., 2022; Creswell and Shanahan, 2022).
Consequences. Higher quality workers may be more expensive (Christoforaki and Ipeirotis, 2014; Huang et al., 2015; Bursztyn et al., 2022). There is also effort and expense involved in filtering crowdworkers for skill (Kulkarni et al., 2014; Zaidan and Callison-Burch, 2011; Valentine et al., 2017) or finetuning models to improve specific skills (Bursztyn et al., 2022).
Examples. Examples that pay particular attention to worker recruitment include Flash Organizations (Valentine et al., 2017), and chains that finetune models (Bursztyn et al., 2022) or use multiple models (Liang et al., 2023).
Related strategies. Worker recruitment works hand in hand with aligning subtasks to worker capabilities. Worker recruitment must also be considered when using adaptive architectures to define what kind of workers to hire during runtime.
5.6. Subtask alignment
To improve outcome quality at each step, align subtasks to worker capabilities (Figure 10).
Title. Subtask alignment, To improve outcome quality at each step, align subtasks to worker capabilities. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Higher quality subtask responses), Constraints (Increased effort to design). Section 2: tactics used. Subtasks (partition, focus), Architectures (dynamic, communicative).
Motivation. One critical strategy is ensuring that subtasks match the workers’ capabilities appropriately. Simplifying tasks can increase outcome quality (Chilton et al., 2013; Wu et al., 2022b; Lasecki et al., 2013; Pietrowicz et al., 2013).
Implementation. Compared to entire tasks, smaller subtasks are often easier to complete, validate, and correct (Ambati et al., 2012; Schick et al., 2022; Wu et al., 2022b; Yang et al., 2022; Kim et al., 2023b; Creswell et al., 2022; Creswell and Shanahan, 2022). Designers can also adapt the subtask content to fit the worker’s capabilities (Chang et al., 2017; Mohanty et al., 2019; Pietrowicz et al., 2013; Arora et al., 2022). For example, researchers found that crowdworkers were better at generating predictive features than at estimating if a feature is predictive; so they adapted the workflow accordingly (Cheng and Bernstein, 2015).
When the appropriate subtasks are unknown, dynamic architectures can use partition or focus subtasks to use workers’ input to break down tasks into an appropriate size at runtime (Kittur et al., 2011; LaToza et al., 2014; Kim et al., 2017; Huang et al., 2023c). Communicative architectures can create appropriate subtasks for workers when they have a running “todo” list from which workers can choose their preferred task of many (Zhang et al., 2012; Lasecki et al., 2015; Mahyar et al., 2018).
Consequences. Much remains unknown about creating appropriate subtasks for LLMs. The capabilities of LLMs and appropriate subtask content are far from fully understood. Furthermore, it is unknown the degree to which LLMs can judge the appropriateness of subtasks for other LLMs to complete, bringing into question the effectiveness of some partition subtask and communicative architecture approaches taken when the workers are crowdworkers. Therefore, testing out these different options to find appropriate subtasks is resource intensive in the design process.
Examples. Examples that pay particular attention to subtask alignment include: Revolt (Chang et al., 2017), CrowdBand (Pietrowicz et al., 2013), and Ask Me Anything (Arora et al., 2022).
Related strategies. Subtasks must align to the capabilities of recruited workers and requirements must be communicated with sufficient context. Subtask alignment is especially challenging when subtasks are created during runtime eith adaptable architectures.
5.7. Context
To promote coherence and improve subtask outcomes, provide the minimum sufficient global context while optimizing for clarity of worker instructions (Figure 11).
Title. Context, To promote coherence and improve subtask outcomes, provide the minimum sufficient global context while optimizing for clarity of worker instructions. Section 1: effects on objectives. Several objectives listed (with their effects). Quality (Critical for quality outputs, Excessive context can confuse), Constraints (Increase effort to design), Accuracy (Affected by small changes). Section 2: tactics used. Subtasks (all), Architectures (communicative).
Motivation. Workflows need to have high quality subtask outputs that promote cohesion among disparate parts of the workflow.
Implementation. Communicating context to workers can help with improving the quality of the output and creating more coherence among disparate subtasks (Luther et al., 2015; Wu et al., 2022a; Teevan et al., 2016; Chilton et al., 2019; Hahn et al., 2016; Ambati et al., 2012; Kulkarni et al., 2012; Bernstein et al., 2010). There are many forms of contextual communication, from input text (Ambati et al., 2012; Kulkarni et al., 2012; Bernstein et al., 2010), to workflow goal tracking (Mirowski et al., 2023; Yang et al., 2022; Kim et al., 2017), to explicit communication between users and workers (Salehi et al., 2017).
Instruction wording is a critical part of all subtask designs to support high outcome quality (Parameswaran et al., 2023). A key difficulty is communicating the user’s intent through language. One study reports that crowdworkers tend to need less explicitly defined instruction than LLMs, but can still struggle with vague or ambiguous instructions (Wu et al., 2023c). Both crowdworkers and LLMs benefit from examples, namely screening and tutorial tasks for crowdworkers (Teevan et al., 2016; Lasecki et al., 2015; Chen et al., 2019; Lasecki et al., 2013; Pietrowicz et al., 2013; Drapeau et al., 2016; Bernstein et al., 2010) and few-shot examples for LLMs (Nair et al., 2023a; Gero et al., 2023). Interfaces for crowdworkers can structure their output, while this structure needs to be communicated to LLMs in natural language (Wu et al., 2023c; Creswell and Shanahan, 2022; Kim et al., 2014).
Consequences. Although context is generally important, overcomplicating instructions can lead to confusion (Chen et al., 2019; Lasecki et al., 2015; Wu et al., 2023c; Madaan et al., 2023b), so understanding the right amount of context to give is important (Kittur et al., 2011). One in-depth crowdsourcing study on providing context through interactions with the user found that providing context was helpful early on in the workflow if content is poor, but became detrimental if communicated later or with high-quality content (Salehi et al., 2017). Minimizing context seen can also help with preserving privacy (Bragg et al., 2021; Kaur et al., 2017). For LLMs, context may be included as part of the prompt, although LLMs have the additional barrier of an input token limit (Park et al., 2022; Mirowski et al., 2023). Therefore, workflow designs should aim for a minimum amount of sufficient context, to provide coherence while avoiding confusion.
For both crowdworkers and LLMs, small changes in instructions or prompts can lead to dramatically different outcomes, impacting the accuracy of the final result (Wu et al., 2022b; Kittur et al., 2012; McDonnell et al., 2016; Wu et al., 2022a; Mirowski et al., 2023; Arora et al., 2022; Huang et al., 2023c; Wu et al., 2023c; Parameswaran et al., 2023; Gero et al., 2023). Additionally, different LLMs may respond differently to the same input prompt (Parameswaran et al., 2023). Therefore, figuring out the clearest instructions requires nontrivial effort and testing.
Examples. Examples that pay particular attention to context include Turkomatic (Kulkarni et al., 2012), collaborative writing with microtasks (Teevan et al., 2016), and a study on communicating context (Salehi et al., 2017).
Related strategies. When using adaptable architectures, dynamically made tasks can struggle to provide quality worker instructions and context. Additionally, the effectiveness of this strategy is limited if there is improper subtask alignment.
Title. Quality thresholds, to more precisely balance the quality-cost tradeoff, measure quality repeatedly and stop the workflow once a quality threshold is reached. Section 1: effects on objectives. Several objectives listed (with their effects). Quality and Constraints (Improves control of tradeoff). Section 2: tactics used. Workers (LLMs, crowdworkers, user), Subtasks (evaluate), Architectures (redundant, dynamic).
5.8. Quality thresholds
To more precisely balance the quality-cost tradeoff, measure quality repeatedly and stop the workflow once a quality threshold is reached (Figure 12).
Motivation. Sometimes, designers may not know prior to running the workflow how much work needs to be done to get a high quality output. In those situations, it would be helpful to have the workflow decide how many resources are necessary to use on an input-specific level.
Implementation. Some workflows give the worker control over the number of redundant generations using evaluate subtasks (Yao et al., 2023; Huffaker et al., 2020; Drapeau et al., 2016; Parameswaran et al., 2023). Others allow for tuning redundant architectures giving flexible goal precision, recall, and agreement thresholds (Huffaker et al., 2020; Yao et al., 2023), use decision-theoretic choices of the number of responses to obtain in redundant workflows (Bragg et al., 2013; Alshaibani and Quinn, 2021; Dai et al., 2010, 2013), or dynamically choose between workflows by their maximum expected value of utility (Goto et al., 2016).
Consequences. Deciding if a threshold is met can require higher costs in terms of effort and expense.
Examples. Examples of this strategy include POMDP-based control (Dai et al., 2013), Pterodactyl (Alshaibani and Quinn, 2021), and a workflow for detecting emotionally manipulative language (Huffaker et al., 2020).
Related strategies. Quality thresholds may be determined by leaning on the adaptive architectures and user guidance strategies.
Figure shows two flow charts, one for crowdsourcing Cascade and one for LLM chain Cascade. Both follow the workflow described in the main text. Several adaptations are highlighted in the LLM chain: the generate step with only one item-category comparison, and the categorize step only looking at the m current items, instead of the entire list.
A line graph showing target number of categories on the X axis from 2-20, and The percentage error on the y axis from -1 to 1, in which 0% error means exactly hitting the target. Two lines are shown. One line, for Cascade-LLM, stays much closer to the 0% error line and has a narrower 95% confidence interval than the second line, zero-shot-target. At 11 or more categories, zero-shot-target has a negative percentage error rate.
6. Case studies
In order to better assess how crowdsourcing methods transfer to LLM chain development, we adapt crowdsourcing workflows to LLM chains in three case studies. This process of adapting workflows supports two investigations. First, it allows us to compare the performance of chaining to both zero-shot prompting and crowdsourcing. Second, through the process of adapting these workflows from crowdsourcing to chaining, we can see how the workflow design changes to accommodate for LLMs. These changes, found through practice, directly compare what tactics support the same strategies and objectives when using crowdworkers or LLMs.
To ensure we are suitably focused, we explore one strategy in more depth: supporting quality control through user guidance in which the user can directly manipulate workflow outputs. We focus on this strategy as it is implemented in crowdsourcing work, but it is underexplored in the LLM chaining space. Prior non-chaining LLM work explores controllability as a helpful attribute (MacNeil et al., 2023; Kim et al., 2023a), so here we seek to incorporate it within chains. With directly manipulable outputs, users can make “rapid, incremental, reversible operations whose impact on the object of interest is immediately visible” (Shneiderman, 1982), increasing control over output elements and helping users explore alternative results (Hutchins et al., 1985). By adapting this subset of the user guidance strategy from crowdsourcing, we explore how and if such a concept can transfer to chains, what modifications are necessary, and its relative advantages compared to controlling the outputs through zero-shot prompting.
We choose three popular crowdsourcing tasks that incorporate different degrees of the accuracy and creativity design space objectives. We implement Cascade (Chilton et al., 2013) to build taxonomies with manipulable structures, Soylent (Bernstein et al., 2010) to shorten text to a variety of target lengths, and Mechanical Novel (Kim et al., 2017) to control the elements of a story. Diagrams of the workflows, as well as our adaptations to use LLMs, can be found in Figures 13, 16, and 18.
For each case study, we first document how the workflow was changed to incorporate LLMs and map those changes to the design space. Common limitations we faced include LLM context lengths, inconsistent output formatting, lack of diversity in model responses, and models struggling to perform the workflow-defined subtasks. We adapted to these limitations with more granular subtasks, more diverse prompts, comparative tasks, and hard-coded validation steps. Across all three case studies, we use different prompts111https://github.com/madeleinegrunde/designingchains at each parallel generation step to increase response diversity, update worker instructions and context through trial and error to get the right output formatting, and structure voting evaluation subtasks to choose among numbered options to minimize hallucinations.
We compare our LLM workflows to zero-shot baselines in terms of output control precision, salient qualities for each task, and the latency and number of API calls. To prioritize human-preferred LLM outputs, we use an instruction-tuned model: OpenAI’s API gpt-3.5-turbo with default parameters (Ouyang et al., 2022). Compared to zero-shot prompting, chained workflows use more resources but can improve output quality. These workflows also provide a greater precision of output control than parameterized zero-shot baselines.
| MIT Indoor Scenes (Quattoni and Torralba, 2009) | MSCOCO (Lin et al., 2014b) | CIFAR100 (Krizhevsky et al., 2009) | ||||
| Zero-shot | Cascade-LLM | Zero-shot | Cascade-LLM | Zero-shot | Cascade-LLM | |
| # of Categories | 2 | 21 | 8 | 24 | 4 | 19 |
| Tree Depth | 1 | 3 | 1 | 4 | 1 | 3 |
| Duplicate Categories | 0 | 0 | 0 | 0 | 0 | 0 |
| Missing Category-Item | 0 | 2 | 1 | 3 | 7 | 2 |
| Incorrect Category-Item | 0 | 0 | 1 | 2 | 2 | 2 |
| Missing Parent-Child | N/A | 0 | N/A | 1 | N/A | 0 |
| Incorrect Parent-Child | N/A | 0 | N/A | 1 | N/A | 1 |
Four taxonomies. The first with default parameters (merge threshold 0.75, minimum category size 2), contains a hierarchy of 4 layers and 23 categories. The second that is more likely to merge (merge threshold 0.5, min category size 2), contains 4 layers and 18 categories. The third that is a larger minimum category size (merge threshold 0.75, min category size 5) contains 3 layers and 13 categories). The fourth that is both more likely to merge and a larger min category size (merge threshold 0.5, min category size 5), contains 3 layers and 9 categories).
6.1. Case study 1: Taxonomy creation
For our first case study, we explore taxonomy creation, the task of organizing a set of items into a category tree to aid sensemaking. This task involves many elements of the accuracy objective, with some degree of creativity: categories need to be correct, but most items can fit into multiple categorizations of different levels of precision. We focus on the ability to control the granularity of these categorizations by controlling the taxonomy shape (Figure 15).
6.1.1. Workflow design
We adapt the Cascade (Chilton et al., 2013) crowdsourcing workflow (Figure 13), which takes as input a list of items and outputs a tree mapping items to categories. The workflow begins by selecting a subset of 32 items, generating category suggestions for each item, evaluating by selecting the best categories among the generated options, and evaluating by confirming which among all categories works for each item. The next step turns the item-category assignments into a taxonomy, deleting small categories and merging similar categories. The workflow then updates the categorizations for the entire list of items with this new assignment. The workflow redundantly iterates with new subsets until all items are categorized, building a growing set of categorizations.
To build Cascade-LLM, our version of Cascade powered by LLMs, we made multiple changes to the original workflow. We highlight some of the most salient modifications. To achieve diverse categorizations, we improved response diversity by giving different prompts to parallel generations and ensuring that every item is input into a category generation step. At first, the categories were too broad. Specifying category qualities via targeted prompts proved unsuccessful, so instead we added a programmatic validation step to delete categories that included “and,” “or,” commas, or slashes. As this hard-coded validation step worked, it was not necessary to include another LLM call. The original workflow gives workers sets of 8 items for which to generate 8 categories. However, the LLM would try categorizing multiple items at once, consistent with prior findings that LLMs struggle to disambiguate multiple subtasks in a prompt (Wu et al., 2023c; Madaan et al., 2023b). To increase diversity, we aligned subtask design to fit model capabilities with one item categorization. Finally, the original workflow also requires parameters for the minimum category size and a merge threshold. To support user guidance through direct manipulation, we dynamically apply these parameters at the last taxonomization stage.
6.1.2. Evaluation setup
We compare Cascade-LLM against zero-shot baselines on three lists of nouns.
Datasets. Our experiments use the labels from the scenes in MIT Indoor Scenes (Quattoni and Torralba, 2009), from object categories in MSCOCO (Lin et al., 2014b), and from CIFAR100 (Krizhevsky et al., 2009). These label sets are of size 67, 80, and 100, respectively. The original Cascade paper explores three datasets sourced from Quora.com question responses. Our LLM workflow is capable of processing these larger datasets, as only individual items need to be within the model’s context size; however, the Quora.com data exceeds the context limit for the zero-shot baselines.
Baselines. We compare our workflow to two baselines. The zero-shot baseline directly prompts the model to create a taxonomy. This baseline establishes how the model completes the task given no restrictions. The zero-shot-target baseline further specifies the target number of categories in the prompt. With this baseline, we compare the ability to control workflow outputs in a single prompt.
6.1.3. Workflows enable direct manipulation
To measure the controllability of the taxonomy’s shape, we measure how well methods can change the taxonomy size (the number of categories). Given a target number of categories (range: 2–20), we measure the percent error between the method’s closest option and the target value.
As seen in Figure 14, one run of Cascade-LLM almost always achieves a precise match with the desired number of categories. In contrast, zero-shot-target is close but usually unable to produce a taxonomy with the target number of categories. The zero-shot condition produces small taxonomies the size of which is not controllable (MIT Indoor Scenes: 2, MSCOCO: 8, CIFAR100: 4).
6.1.4. Workflows reduce hallucinated and forgotten items
We also evaluate the quality of the outputs through qualitative evaluation. Similar to the original paper, we hand-code taxonomies without knowledge of which method created it. First, we look for errors in items across all baselines. Specifically, we see if the output hallucinates or forgets any items. We then code the zero-shot and default-parameter Cascade-LLM baselines for the same mistakes considered in the original paper: duplicate categories, missing parent-child relationships, and incorrect parent-child relationships.
Cascade-LLM does not forget any items, but 83.72% of zero-shot taxonomies do not include all items. Cascade-LLM cannot hallucinate new items, while zero-shot hallucinates 51 items across the 129 taxonomies. Cascade-LLM taxonomies have similar error rates to zero-shot taxonomies on duplicate categories and missing/incorrect category-item and parent-child pairs, yet are richer taxonomies with more categories and depth (Table 1).
6.1.5. Workflows have higher costs
We measure generation time and the number of API calls to an LLM. Cascade-LLM takes more time (mean: 2,642.58 sec., std: 655.81) than the zero-shot (mean: 24.35 sec., std: 5.04) and zero-shot-target (mean: 21.69 sec., sd: 4.49) baselines. Crowdsourced Cascade took 43 hours and 3 minutes to taxonomize three item sets (Chilton et al., 2013). Cascade-LLM makes more model calls (mean: 18,898.33, std: 4472.61) than the one call from the zero-shot and zero-shot-target baselines. Most of these calls are short in terms of token length, as they compare one item to one category. Crowdsourced Cascade reports 1760 calls to workers per iteration. For our goals, the increased cost in time and computation enables direct manipulation on more detailed categorizations without item forgetting or hallucination.
Figure shows two flow charts, one for crowdsourcing Soylent and one for LLM chain Soylent. Both follow the workflow described in the main text. Several adaptations are highlighted in the LLM chain: 1) the initial text splits into smaller chunks, 2) the Find and Fix steps are split into parallel edit, merge, and delete steps, and 3) automated format validation is added after the fix step.
6.2. Case study 2: Text shortening
The next case study is the task of shortening text. Unlike summarization, shortening the text involves making edits to create a condensed version of the text that attempts to preserve the full meaning. Shortening is a task on the creative side, but requiring elements of accuracy. Shortened versions must stay faithful to the original text, and there are many options of what phrases to shorten and how to shorten them. By implementing a workflow to generate many shortening options, an end user can control the final length of a paragraph dynamically.
6.2.1. Workflow design
To shorten text we adapt the Find-Fix-Verify workflow (Figure 16) introduced in Soylent (Bernstein et al., 2010). In the “Find” step, crowdworkers select areas that might be shortened. If at least two workers agree on a text span, that span progresses to the “Fix” step. In the “Fix” step, crowdworkers make edits to shorten the text. In the final “Verify” step, more crowdworkers verify the possible changes by rejecting by majority vote any edits deemed grammatically incorrect or that change the meaning of the sentence. These steps respectively map to the focus, transform, and evaluate subtask types in our design space. The workflow outputs edit options for multiple spans in the paragraph, which in turn may have multiple possible edits. Soylent defaults to showing the shortest possible option, but it also provides a length slider for the user to adjust the final output as desired.
We made multiple adjustments to produce our Soylent-LLM workflow. First, we ran into context limits and parallelization constraints with longer texts. In response, we apply a dynamic approach that splits the workflow into chunks of 10 sentences and applies the workflow to each. Second, Soylent allows workers to fix inputs by editing, merging, and deleting phrases. LLMs performed better when we split these tasks into separate prompts, so we further decompose Find and Fix into finer grained editing, merging, and deleting steps to get an align subtask size to the model’s capabilities. Finally, in the Find step the LLM would sometimes “hallucinate” phrases to shorten that do not exist in the text or would change quotes. To mitigate these errors, we add several hard-coded validation steps: we ensure that the phrase to replace exists in the input text, that the replacement is shorter, and that no quotes were altered or added.
Recent work also implemented three unique variations of a Find-Fix-Verify workflow using LLM chains (Wu et al., 2023c). Their evaluation, using student preferences, found mixed results on effectiveness for creating better shortenings than a zero-shot baseline. We evaluate our workflow along different dimensions: controllability of output length, fluency, grammaticality, what type of content is deleted, and workflow resource usage. Unlike this prior work, we also retain the parallel generation and validation steps from the original paper.
6.2.2. Evaluation setup
For evaluating Soylent-LLM, we use texts from newspapers and from the original Soylent paper (Bernstein et al., 2010). We compare Soylent-LLM with three zero-shot baselines.
Dataset. We use the XSUM summarization dataset (Narayan et al., 2018). We only use the input texts because our task is shortening, not matching gold-standard summarizations. We use 100 randomly sampled texts from XSUM, with the requirement that examples are more than 500 words long. For qualitative analysis, we use the texts from the Soylent evaluation (Bernstein et al., 2010). Through this dataset, we can compare the Soylent-LLM shortenings to Crowdsourced Soylent shortenings.
Baselines. The zero-shot baseline prompts the model to shorten the paragraph directly. For the zero-shot-target baseline, we specify the target output length in a zero-shot prompt. In the zero-shot-ffv baseline, we ask the LLM to explicitly perform Find-Fix-Verify steps in one prompt.
Two graphs side by side. The first graph is a line graph graphing the target number of words on the X axis from 100-500 and percent error on the y axis from -75% error to 250% error. The graph contains 4 lines. On short texts, the zero-shot-target gets the lowest error followed by zero-shot, Soylent-LLM and zero-shot-ffv. On longer texts Soylent-LLM has approximately 0% error. Zero-shot has the highest error, and zero-shot-target and zero-shot-ffv have approximately the same error at 50%/ The 95% confidence intervals are narrow. The second graph is a scatter plot with the X axis the target percent length of original ranging from 0 to 100 and with the Y axis the percentage error from -100 to 500%. There are 4 colors of points. The highest percent errors are for small target lengths across all 4 baselines. As the target percentage increases, Soylent-LLM converges to close to zero percent error, while the three zero-shot baselines trend negative.
6.2.3. Workflows enable length manipulation
We first compare how well the various LLM approaches reach a desired output length. We represent target lengths through word counts of 100-500 words. In initial experiments, we found that zero-shot baselines perform more competitively when asked to shorten to a given word count rather than to a length percentage and hence use word counts in targeted prompts. Nevertheless, we also measure accuracy in terms of a target percentage of length. We calculate the target percent length by dividing the target number of words by the input text’s number of words, and we calculate the actual percent length as the output text’s number of words divided by the input text’s number of words. We compare the percent difference (i.e., error) between the actual and target percent lengths.
We find that as text length increases, Soylent-LLM can more precisely meet the target length (Figure 17). When shortening text, the zero-shot prompt condition consistently outputs approximately 100 word long shortenings (mean: 126.14, std: 52.63). Thus, for longer target lengths, this method achieves lower performance. Zero-shot-target is similarly best at lengths close to 100 words, but it still shows a lack of precision in reaching longer target lengths. Soylent-LLM and zero-shot-ffv both struggle to hit shorter targets. This result matches the Soylent paper, in which output texts were between 78-90% of the original text length (Bernstein et al., 2010). Soylent-LLM can more precisely hit the larger target lengths than any of the baselines.
We find similar trends measuring percent error with the percent length reduction required to meet the target length. Soylent-LLM and zero-shot-ffv can perform poorly with large reductions, in which the target % is a small fraction of the input text, while zero-shot and zero-shot-target perform worse with smaller reductions. In summary, Soylent-LLM does a better job of tightening long texts but is less well-suited to extreme shortenings that cross over into summarization. In addition, Soylent-LLM is unique in providing options among all targets without having to requery. Moreover, at each target length, there are often multiple possible output texts.
6.2.4. Workflow outputs have abrupt transitions but lower perplexity
We measure the average perplexity of output texts as a proxy for fluency. We use a different language model (Facebook’s OPT-1.3b) than the worker model (OpenAI’s gpt-3.5-turbo) for an external evaluation. We also perform a qualitative analysis of the shortening procedure. We use the five texts used for evaluation in the original Soylent paper (Bernstein et al., 2010). We compare the edits made with the shortest output from Soylent-LLM, the baseline conditions, and the crowdsourced condition’s output provided by the authors of the original paper. Without knowing the condition, we coded the outputs for fluency and grammatical errors, following the qualitative evaluation procedure of the original paper. We additionally coded what content changed during shortening.
Soylent-LLM has the better result of a lower perplexity (mean: 11.69, std: 3.77) than zero-shot (mean: 15.37, std: 5.73), zero-shot-target (mean: 12.31, std: 3.91), and zero-shot-ffv (mean: 16.19, std: 5.68). This result surprised us, as zero-shot is fully generated by a large language model fine-tuned to provide fluent responses.
As with the original crowdsourcing workflow, some shortenings diminished fluency or introduced grammatical errors. Zero-shot, zero-shot-ffv and Soylent-LLM shortenings all included passive voice, but were otherwise grammatical. However, Soylent-LLM and zero-shot-ffv both could contain abrupt endings or transitions. This abruptness is due to separately editing parts of the input text. Crowdsourced Soylent induced grammatical errors as well as abrupt transitions. We found that zero-shot and Soylent-LLM both shorten by removing unnecessary phrasing and deleting examples and justifications. Zero-shot more successfully merges sentences than Soylent-LLM, while Soylent-LLM changes less text and more successfully retains the user’s voice, rather than formalizing it. Soylent-LLM’s changes affect a larger portion of the text than Crowdsourced Soylent. In sum, Crowdsourced Soylent made 42 edits to phrases and 12 to entire sentences, while Soylent-LLM made 12 edits to phrases and 36 to entire sentences.
Figure shows two flow charts, one for crowdsourcing Mechanical Novel and one for LLM chain Mechanical Novel. Both follow the workflow described in the main text. Several adaptations are highlighted in the LLM chain: 1) additional validation steps in the reflect part of the workflow, 2) an automated length validation step in the revise part of the workflow 3) extra iterations through the scenes of the text as well as with different seeded texts.
6.2.5. Workflows have higher costs than zero-shot prompting
We measure the time it takes to generate the outputs, as well as the number of API calls to the model. The number of API calls varies depending on the length of the input text and the number of items that are selected at each step.
Soylent-LLM takes much longer (mean: 105.44 sec., std: 60.16) than the zero-shot (mean: 9.09 sec., std: 2.92), zero-shot-target (mean: 13.65 sec., std: 5.47), and zero-shot-ffv (mean: 43.55 sec., std: 23.16) baselines. Crowdsourced Soylent takes longest with a reported range of 2,640–29,340 sec. (44-489 min.) (Bernstein et al., 2010). Compared to the one call to the model for the zero-shot baselines, one Soylent-LLM run averages 127.11 calls to the model (std. 63.56). Crowdsourced Soylent reports incorporating 158–362 workers.
6.3. Case study 3: Short story generation
The final case study concerns story generation given a short text prompt. This task lies on the open-ended creativity side of the spectrum, as short stories have a wide design space of possible plots, characters, messages, imagery, and lessons. In this task, user guidance of the story involves higher level control of different plot elements, with the ability to see counterfactuals (e.g., if a story element is included or not). We implement a workflow that, compared to zero-shot approaches, enables this exploration more precisely, while also creating stories that are preferred by crowdsourced raters for imagery, originality, and style.
6.3.1. Workflow design
We adapt the Mechanical Novel (MN) (Kim et al., 2017) workflow for this task. This workflow involves three main steps (Figure 18). The first generate subtask “initializes” a story with multiple scenes based on a given story prompt (e.g., “Malcolm finds himself alone in a runaway hot air balloon and accidentally travels to a city in the sky.”). Then, the “reflect” and “revise” steps are iterated five times in a loop. In the “reflect” step, workers consider what changes might helpfully adjust the story. In the “revise” step, workers implement those changes. Each step has subtasks for parallel generation of suggestions and edits, as well as evaluate subtasks voting for the best generation.
We adapted the Crowdsourced MN to MN-LLM by incorporating fine-grained user guidance with manipulatable outputs. At the conclusion of the workflow, instead of receiving just one or more separate short stories, the user can interact with different combinations of edits to compare stories with different quality combinations and compare counterfactuals on editing decisions. To enable these outputs, we run the “revise” stage on texts of all combinations of suggestions from the “reflect” stage.
Beyond introducing control of story elements, the largest challenge was managing token context limitations. Getting response diversity with iterative improvement on texts often increases the text length (Little et al., 2010a). To account for context length, we generate stories with four scenes rather than six and iterate through scenes rather than informing workers to select scenes. In our design space this change maps to using a redundant architecture rather than a sequential architecture with a focus subtask. At the beginning of the “revise” step, we add a generation subtask that summarizes the current story to provide abbreviated context in the worker instructions. We included a hard-coded validation step with an evaluate subtask to disqualify edits that were more than 1.5 times the input text length. Unfortunately, this disqualified many highly descriptive texts and sometimes meant that no edits were made because all suggestions were too long.
| Imagery | Coherency | Plot | Originality | Style | Technical | Overall | |
| MN-LLM | 124 | 97 | 93 | 137 | 118 | 97 | 112 |
| Zero-shot | 76 | 103 | 107 | 63 | 82 | 103 | 88 |
| Zero-shot | MN-LLM | |
| Avg. # combinations | 3.80 | 20.40 |
| Avg. length in words | 276.38 | 642.93 |
| Precision | 77.05 | 97.74 |
| Recall | 88.68 | 90.00 |
| Suggestion type | Zero-shot | MN-LLM | Passed | Voted |
| Add plot events | 25 | 30 | 13 | 6 |
| Introduce character | 14 | 18 | 13 | 3 |
| Introduce significant object | 6 | 1 | 1 | 0 |
| Add plot twist | 4 | 11 | 10 | 3 |
| Add detail to interactions | 1 | 6 | 3 | 1 |
| Add emotional detail | 0 | 17 | 8 | 3 |
| Add background details | 0 | 8 | 5 | 0 |
| Add a different perspective | 0 | 5 | 2 | 2 |
| Add a moral lesson | 0 | 2 | 1 | 1 |
| Add story detail | 0 | 2 | 1 | 0 |
Our initial pilots generated editing suggestions that were very generic (e.g., “Add more detail to the story”) and were repetitive across iteration rounds. To improve the diversity of suggestion outputs, we tried changing the worker instructions to be contrastive by providing previously used “banned” suggestions. However, the outputs anecdotally seemed more likely to follow those of prior rounds after this change. We hypothesize that the length of the context and request to do two tasks in one (generate a suggestion and avoid the prior suggestions) led to a lack of diversity. Instead, we add several validation steps to promote suggestion diversity and quality. This subtask granularity worked better. In these evaluate subtasks, we prompted the LLM to ensure that 1) the suggestion did not overlap with prior suggestions or two seeded generic suggestions (“Generic recommendation to introduce conflict.” and “Generic recommendation to add detail.”), 2) the suggestion has not already been implemented in the story, and 3) that the suggestion is worded as a suggestion, not new text for the story.
6.3.2. Evaluation setup
To evaluate MN-LLM, we use the story prompts and evaluation procedure from the original paper (Kim et al., 2017) and compare MN-LLM’s results to those of two baselines.
Dataset. Our dataset consists of the five short story prompts in the Mechanical Novel paper (Kim et al., 2017).
Baselines. The zero-shot baseline method prompts the model to generate five suggestions, then write one story that incorporates these suggestions. The zero-shot-combo baseline method prompts the model to generate five suggestions, then write different versions of a story based on combinations of those suggestions.
6.3.3. Workflows enable control of plot elements
We assess each output to evaluate the number and quality of suggestion combinations that emerge from MN-LLM and zero-shot-combo. To measure quality, we calculate the length of texts, as well as the precision and recall of the included suggestions. To measure precision, we calculate what percentage of included suggestions were supposed to be included. To measure recall, we calculate what percentage of suggestions that were supposed to be included were included. As MN-LLM is iterative, some later suggestions reference elements initialized from earlier suggestions, so when coding, we consider the parts of each suggestion that differ from one another.
As shown in Table 3, MN-LLM generates more, longer, and higher quality suggestion combinations than the zero-shot baseline. MN-LLM does not always generate the 33 story options afforded by the initial story and combinations of 5 suggestions, as sometimes no generated suggestions passed our validation steps, or all generated fixes were too long. However, it still generates more than 20 options on average, in comparison to zero-shot’s average of 3.8.
6.3.4. Workflows generate a wider variety of suggestions
Without knowing the condition, we coded the story suggestions in the MN-LLM and both zero-shot conditions by theme. MN-LLM suggestions were more varied in type than the zero-shot suggestions (Table 4). The MN-LLM validation steps removed nearly half of the generated suggestions and most often voted for adding plot, characters, or emotional details as the final suggestions.
6.3.5. The MN-LLM workflow improves imagery, originality, and style
Following the evaluation methods of the MN paper, for each story prompt we gave the output from MN-LLM and the zero-shot baseline, in randomized order, to 200 workers on Amazon Mechanical Turk using EasyTurk (Krishna, 2019). We asked crowdworkers to vote between stories from the Soylent-LLM and zero-shot baselines along seven story writing dimensions (Kim et al., 2017). We paid workers $1.50 USD, the equivalent of $18 per hour.
The Mechanical Novel workflow improved the quality of the output narratives, although sometimes at the expense of a concise plot. We find that MN-LLM stories were more highly rated along the Imagery, Originality, and Style, and Overall dimensions. Zero-shot-generated stories gained more votes in the Coherency, Plot, and Technical dimensions (Table 2).
6.3.6. Workflows can struggle with coherence
Without knowing the condition, we also reviewed the MN-LLM and zero-shot conditions for plot irregularities. Long-form plot coherence of creative LLM outputs is difficult to achieve (Mirowski et al., 2023; Yang et al., 2022), and we find some plot irregularities in the MN-LLM outputs. Three stories revealed the same information multiple times. Two plots had disjointed elements, one in which the same pronoun “she” referred to two characters in confusing ways and one with a time and location skip due to a scene that was never selected for editing. Such plot irregularities occur in the MN-LLM outputs because of the iteration through scenes and the use of a story summary as context. Both of these design choices were made to avoid exceeding LLM context limits.
6.3.7. Workflows have higher costs
We measure the time and number of API calls required to run the workflow compared to running the baselines. Compared to zero-shot (mean: 21.88 sec., std: 5.89), MN-LLM takes much longer in seconds to complete (mean: 1,452.97 sec., std: 930.26). Compared to zero-shot’s one call, MN-LLM makes 2,398 LLM calls on average. The crowdsourced Mechanical Novel paper does not report latency or number of workers (Kim et al., 2017).
7. Discussion
The approach of decomposing a task into a series of subtasks is conceptually the same for both crowdsourcing workflows and LLM chains. Through overlap in the literature and our case study observations, we find that the same objectives and tactics apply to both, with the exception of the choice of workers. We also find that the strategies found in crowdsourcing apply to chaining, as do their impact on objectives. Despite these similarities, LLMs and people have different abilities, strengths, and weaknesses, so we sometimes need different approaches for employing tactics to effectively implement the same strategies.
Thus, strategies and their effect on objectives transfer directly from crowdsourcing to LLMs. However, the tactic designs to support strategies can vary significantly, and could benefit from improved development and testing processes. Understanding how to effectively use tactics to implement strategies is not only a more focused goal than the connection between tactics and objectives, but it is also more generalizeable. For example, learning how to effectively get LLMs to validate their own outputs would be applicable to many different tasks and objectives.
In this section, we discuss further the similarities and differences between LLM chains and crowdsourcing workflows, lessons that exploring this intersection has taught us about using tactics to support strategies in LLM chain design, and unknowns that mark promising avenues for future discovery.
7.1. LLM stochasticity is different than human diversity
One of the key benefits of crowdsourcing workflows is that they incorporate multiple people, such that crowdworkers can fill in each other’s deficits (Huang et al., 2017; Nebeling et al., 2016; Kim and Monroy-Hernandez, 2016; Zhang et al., 2012). Therefore, requesting multiple crowdworkers for the same or sequential subtasks inherently leads to differing outputs and considerations. The diversity of perspectives happens inherently by querying different workers but is limited to available crowdworkers and thus affected by demographic and geographic disparities on crowdsourcing platforms (Difallah et al., 2018).
LLMs also have the capability to output different responses, as their training data includes information produced by many people on the Internet (Bender et al., 2021). To induce diversity, prior work has tried methods like adjusting prompts, temperature, and model types (Parameswaran et al., 2023; Park et al., 2022; Du et al., 2023; Arora et al., 2022). However, ongoing work has found that there are elements of diversity in people that are not evident in LLMs. For example, LLMs struggle to output pluralistic responses (Sorensen et al., 2024), are limited to the outputs of people with access to the Internet (Bender et al., 2021), and cannot represent identity groups (Wang et al., 2024). Our case studies suggest that LLMs need more variation in prompting and explicit intervention to get diverse responses, although we did not test this systematically. Intuitively, this makes sense because there is great diversity in the human population, but models are known to mimic the majority voice to minimize their objectives, ignoring marginalized opinions that make up a smaller percentage of the training data (Gordon et al., 2022).
Using a high temperature and prompting the LLM with a variety of different instructions can improve diversity somewhat but is unlikely to match that of crowds. Although this difference between workers may at some level seem obvious, it impacts the tactics needed to effectively use diverse responses. Diverse responses are used both to source variety and to check agreement. In crowdsourcing, the underlying tactic behind both of these goals is identical: run the same subtask with multiple workers. However, different tactics can better explicitly address these applications for LLMs.
Variety. Diverse responses are often used to support creative tasks with a variety of outputs. Consider one step in Mechanical Novel that generates suggestions for updating the story. As this is an iterative workflow, we want a variety of story update suggestions both at each iteration and across iterations. Unfortunately, even with variety in generation prompt wording, we found a low diversity of suggestion content (Table 4 “Zero-shot” column). After repeated challenges in prompt engineering to increase the variety of suggestions, we eventually intervened with an added validation step to ensure diversity of story updates. This approach was effective, but inefficient, as it filters out many generated responses. Role-based communicative architectures are also promising in giving prompts extra power to increase variety. Although giving identity based roles should treated with caution (Wang et al., 2024), giving other roles to LLMs could be a more powerful way of inducing a variety of responses (Park et al., 2022). Many roles in crowdsourcing workflows were based on organizational positions for relevant domains such as journalism (Agapie et al., 2015) or project development (Valentine et al., 2017).
To source variety, the best choice of redundant architecture can differ when using crowdworkers or LLMs. Prior crowdsourcing work has found that parallel and iterative architectures source different degrees of diversity and quality of responses in crowdsourcing depending on the task (Goto et al., 2016; Little et al., 2010a; Kittur et al., 2011; André et al., 2014). For example, iterative architectures have less variation and a higher average quality of outputs, but have a lower maximum quality output (Little et al., 2010a). Some of these tradeoffs will hold with LLMs, but differences may emerge as well. For instance, some work suggests that, relative to crowdworkers who experience anchoring bias on iterative architectures, LLMs can perform divergent thinking to generate something explicitly different than provided examples (Wu et al., 2023c, a). Similar studies on LLM chains would better inform the effect of different redundant architectures on variety and quality.
Agreement. Crowdsourcing work has used agreement among crowdworkers as a proxy for confidence vs uncertainty (Liem et al., 2011; Chang et al., 2017). This agreement can be used to validate the quality of outputs, so understanding how agreement relates to quality is of critical importance. Here, we focus on agreement among outputs of identical tasks, but the difference between a different person validating the outputs of another, and a model sequentially validating its own outputs, should also be considered.
There are many potential methods for measuring confidence for LLMs, including a confidence score, checking consistency across small prompt variations, and querying multiple times for taking the majority answer (Xie et al., 2023; Parameswaran et al., 2023; Liang et al., 2023; Du et al., 2023). The effect of each of these approaches on the quality of answers is not well understood (Xie et al., 2023; Parameswaran et al., 2023; Liang et al., 2023; Du et al., 2023). It is unclear the degree to which just the consistency in one LLM’s outputs is a proxy for confidence and accuracy, versus consistency across outputs from varied prompts or models.
This more tenuous connection between agreement and correctness may have some impact on chosen tactics. For example, communicative architectures in crowdsourcing were often based on the premise that agreement among people mattered, but that may not transfer to LLMs. Consider debate-based communicative architectures. Some crowdsourcing studies only bring up debate if there’s disagreement (Chen et al., 2019; Drapeau et al., 2016; Chang et al., 2017), but LLMs may benefit more from inducing debate for all inputs, rather than if two outputs disagree.
Therefore, the differences between human diversity and LLM stochasticity impact the design of redundant and communicative architectures to support divergent thinking or agreement.
7.2. LLMs validate response quality less reliably than crowdworkers
Validation in crowdsourcing workflows often involves asking a different worker to verify if errors or desirable qualities exist in outputs. However, both the literature and our case studies show mixed results on the effectiveness of this strategy for chaining.
While some studies show that prompting a model to reflect can help achieve higher accuracy outputs (Madaan et al., 2023b; Xie et al., 2023; Gero et al., 2023; Yao et al., 2023), others caution against the same model validating its own outputs, as LLMs exhibit high confidence for incorrect prior outputs (Liang et al., 2023; Du et al., 2023). Some approaches for more robust validation include: validating with a different LLM (Liang et al., 2023), debate-based communicative structures (Liang et al., 2023; Du et al., 2023), and structural validation such as similarity across parallel redundant streams (Liem et al., 2011). One paper found that “ease of persuasion,” or how quickly an LLM relented in a debate, was more indicative of factual accuracy than other measures of confidence (Du et al., 2023).
Through our case studies, we found several helpful practices for adapting validation steps to work with LLMs.
First, validating “quality” is open-ended, so it is helpful to specify aspects of quality to check instead of using one general “quality” validation step. Crowdsourced Soylent already split the validation into separately checking for “grammar” and for “meaning” changes (Bernstein et al., 2010). However, we needed to make these checks even more targeted to adapt to LLM-specific flaws. Crowdsourced Soylent did not need to check that the rewritten text was shorter, but we added a length-checking step to Soylent-LLM after struggling to ensure a shorter length through prompting alone. We also explicitly validated that the model did not change direct quotes, a nuance that is understood in crowdworkers’ common sense. However, the subphrases from Soylent-LLM were for the most part fluent, deprioritizing the importance of the grammaticality check. In another example, crowdsourced Mechanical Novel involves one evaluate step in which crowdworkers choose the “best” suggestion. We added three validation checks to MN-LLM before this step to account for common errors in chosen suggestions. We explicitly check that LLM responses 1) do not overlap with prior suggestions, 2) are not already implemented in the story, and 3) are worded as a suggestion rather than as new text. Although we did not specify any in our case studies, LLMs may require additional validation of output structures, as text-based responses cannot be restricted by interface design the way crowdworker interfaces can.
Validation should address specific elements of quality that are aligned to worker capabilities. In our case studies, we determined what elements of quality to validate by evaluating intermediate outputs in a trial and error process. Determining these qualities can be tricky, especially as the capabilities of LLMs rapidly change. To determine what elements of quality require validation, intuition needs to be tempered by empirical testing and subtask evaluation.
Even after the set of qualities under consideration has been determined, we found LLMs could still struggle with successfully validating these qualities. To adapt to these issues, we used several approaches. First, when possible we made evaluate steps programmatic rather than LLM-based. This approach is most reliable, but there are a limited number of domains where programmatic checks are feasible. Second, we continued to break qualities down into even simpler parts when needed to align subtask difficulty with worker capabilities. Third, we simplify context to the minimum needed. Considering the story suggestion quality checks above, we found that including multiple prior suggestions, or including the entire story when not necessary, reduced the quality of validation checks.
Some types of validation remain difficult for LLMs. Anecdotally, we found that some more abstract validation tasks, such as identifying changes in meaning, were challenging to enact with LLMs. In Crowdsourced Soylent, the “Verify” step evaluates if the fix changes the meaning of the statement. This requires nuance for what elements of the meaning are necessary. Overall, supporting specificity and simplicity in validation steps were key for adapting these workflows to use LLMs.
7.3. Chains for subjective tasks can use crowdsourcing approaches to enable users to navigate the space of outputs
Some tasks have a large space of possible correct outputs, due to uncertainty or creativity, and could benefit from the user giving direct input. Under these conditions, it is helpful to draw on longstanding work on mixed-initiative user interfaces (Horvitz, 1999) to incorporate users into the workflow. Such an approach allows for parts of the workflow to be automated while others involve user input.
By enabling user initiative on workflow direction or outputs, users can navigate the space of outputs—either to explicitly take control or to learn through the process of exploration. For example, this could be helpful for determining classification label sets. Instead of predefining all possible rules in instructions, workflows can surface uncertain areas for users to make final decisions (Chang et al., 2017). In more creative ventures like prototyping, incorporating user feedback to align the output to their vision can avoid technically correct but irrelevant outputs (Kulkarni et al., 2014).
User control through prompting is brittle (Arora et al., 2022), so workflows can enable more fine grained control. LLM chains can draw from user interaction methods from crowdsourcing workflows. Several large types of user interaction emerge from the literature.
First, the user may act as a step in the workflow for validation or for clarifying context when triggered by ties (Pietrowicz et al., 2013) or crowdworkers (Kulkarni et al., 2012; Salehi et al., 2017). The main design decisions for this approach are deciding what and when the user will validate. The structure and timing of user interaction can matter, for example when supplying workers with needed context (Salehi et al., 2017).
Second, the user and crowdworkers communicate through a shared interface in a relatively unstructured cycle of interactions while creating a shared creative output. Hurdles here include making user review of outputs easier and reducing latency (Nebeling et al., 2016; Lasecki et al., 2015). Metaphorian is a chaining paper that follows a structured version of this in which users work on an interface to develop metaphors, but are limited to a few predefined actions (Kim et al., 2023b).
Third, a user may review options between iterations of the workflow (Willett et al., 2012; Cheng and Bernstein, 2015; Kulkarni et al., 2014; Park et al., 2022; Mirowski et al., 2023). This approach can help with exploring design spaces (Park et al., 2022; Kulkarni et al., 2014; Liu et al., 2023) and determining when the final quality threshold is met (Willett et al., 2012).
Fourth, a user may directly manipulate the output to see the breadth of user responses and to make final decisions about the output with less effort than doing the task directly (Chang et al., 2017; Chilton et al., 2013; Willett et al., 2012; Cheng and Bernstein, 2015; Lee et al., 2017; Mohanty et al., 2019; Bernstein et al., 2010). This approach enables users to surface uncertainty in the data and express control over the final output.
Many chaining papers discuss a generic possibility of a user replacing steps or intervening if needed (Schick et al., 2022; Yang et al., 2022; Wu et al., 2023a; Wang et al., 2022; Wu et al., 2023b; Zhang et al., 2024), but few actually implement this with people (Park et al., 2022; Kim et al., 2023b; Mirowski et al., 2023). We can see from crowdsourcing that there are many ways of incorporating the user that bring myriad benefits. In our case studies, we take direct manipulation from crowdsourcing and apply it to chaining, increasing the controllability of LLM outputs.
7.4. On-demand availability of LLMs reduces barriers to real-time crowdsourcing architectures
Unlike crowdworkers, LLMs are available to be queried at all times with low latency. While, in one sense, this observation is obvious, it has deep implications on tactics.
The unpredictable availability of crowdworkers presented challenges for workflows in which crowdworkers must stay available for extended periods of time or must be called on-demand (Bernstein et al., 2011). While incentive schemes could provide some real-time response, incentive tuning and implementation added significant complexity (Lasecki et al., 2013; LaToza et al., 2014; Weir et al., 2015). As a result, several crowdsourced workflows — while achieving impressive results on paper — were too complex for widespread deployment (Chen et al., 2019; Valentine et al., 2017). The availability of LLMs reduces the barriers to these workflow designs and their benefits.
Methods like retainer models that incentivize keeping crowds were important for crowdsourcing workflows that involve real-time interaction for users (Bernstein et al., 2011) and for communicative workflows like debate in which workers must remain online as other workers respond to them (Drapeau et al., 2016; Chen et al., 2019). These workflows improve latency for users and improve quality and accuracy, respectively, but the challenges of retaining workers increased barriers in crowdsourcing. Similarly, for mixed-initiative interfaces, delayed worker responses increased the latency of interactions (Lasecki et al., 2015). Reduced latency of worker responses can make user interactions more real-time.
In other workflows, such as Flash Organizations (Valentine et al., 2017), workers recruit other workers to complete subtasks that are dynamically created. Finding the right workers for the job took a median of 14 minutes with a light screening process and of 15 hours with a more intense screening process (Valentine et al., 2017). In contrast, the Orchestrator LLM of the AutoGen (Wu et al., 2023a) chain “recruits” with other calls to LLMs, which can happen flexibly and quickly (Wu et al., 2023a).
Therefore, using LLMs as workers reduces barriers to using powerful communicative workflow regimes.
7.5. LLMs need different, often simpler, subtasks
Throughout our case studies, we often needed to break crowdsourcing subtasks into multiple, more specified subtasks. Multiple factors necessitated these changes: natural language is difficult to structure (e.g., thus, we query for one item category at a time rather than eight in Cascade), LLMs have difficulty balancing multiple requirements in a subtask (Wu et al., 2023c; Madaan et al., 2023b; Hsieh et al., 2023) (e.g., in response, we split Soylent’s “Fix” step into separate edit, merge, and delete queries), and LLMs do not understand length-related constraints (e.g., to address this, we add a hard-coded validation step after a generate subtask in Mechanical Novel to check for length).
In our case studies, LLMs performed better on highly specified subtasks for chains rather than fewer, more complicated subtasks. However, this may change as newer, more powerful LLMs are introduced.
7.6. Instructions should provide the minimum necessary context
Providing global context to subtasks is widely agreed to be a critical element of subtask design and output coherence, but it can increase instruction complexity (Wu et al., 2022a; Salehi et al., 2017; Ambati et al., 2012; Luther et al., 2015). However, findings from the literature survey and case studies suggest that LLMs can respond poorly to long, complicated instructions (Du et al., 2023; Parameswaran et al., 2023; Madaan et al., 2023b; Wu et al., 2023c). This context limitation should be re-evaluated over time as LLM context windows grow.
The nuances of context inclusion appeared in our case studies. For example, we wanted to provide MN-LLM the entire short story thus far as context, as occurs in crowdsourced Mechanical Novel. To generate diverse suggestions for improvement we gave the LLM previous suggestions as context for what it should avoid. However, adding this requirement to the overall prompt anecdotally increased the output’s chance of anchoring to prior suggestions. Therefore, we made a separate evaluate subtask that determined if the new suggestion matched previous suggestions out of context from the rest of the story.
We recommend designers take into consideration that LLMs may not adequately process all the nuances of complicated instructions. Simplifying the subtask size in turn simplifies instructions and, often, the amount of necessary context.
8. Future opportunities for chaining
In this section, we highlight practices in crowdsourcing that we consider likely to be worthwhile to transfer to chaining. We encourage future work on these topics.
8.1. Creative tasks with dynamic and communicative architectures
Crowdsourcing studies currently support a wider range of tasks for creative work with a wider range of approaches than chaining studies. We expect these approaches to transfer to chaining and bring myriad benefits.
Future work could attempt creative tasks that are under-explored in chaining, such as: prototyping (Lee et al., 2017; Valentine et al., 2017; Kulkarni et al., 2014), categorizing (Huffaker et al., 2020; Bragg et al., 2013; Chang et al., 2017), topic concept maps (Liu et al., 2018; Hahn et al., 2016; Luther et al., 2015), project development (Mahyar et al., 2018), or more specific tasks like itinerary creation (Zhang et al., 2012), researching a purchase (Kittur et al., 2011), and conceptual blends (Chilton et al., 2019).
Crowdsourcing work also uses a broader range of communicative and dynamic architectures to support creative work. Giving workers roles during communication can assist both crowdsourcing workflows (Valentine et al., 2017; Agapie et al., 2015; Retelny et al., 2014) and LLM chains (Wu et al., 2023a). In crowdsourcing work, some studies go beyond giving a role in instructions and instead create free-form role-based (Nebeling et al., 2016; Valentine et al., 2017; Zhang et al., 2023b), hierarchical (Agapie et al., 2015; Retelny et al., 2014), and state-sharing (Lasecki et al., 2015; Lee et al., 2017; Nebeling et al., 2016; Kulkarni et al., 2014; Mahyar et al., 2018; Lasecki et al., 2013) communicative architectures.
Dynamic architectures provide more flexibility when contingencies arise (Kim and Monroy-Hernandez, 2016; Valentine et al., 2017; Kittur et al., 2011; Kulkarni et al., 2012) and adapt to edge cases (Kim et al., 2014; Huang et al., 2015; Park et al., 2022; Kim et al., 2023b; Schick et al., 2022; Creswell and Shanahan, 2022; Yao et al., 2023; Zhou et al., 2022b; Nair et al., 2023b; Kulkarni et al., 2012; Hahn et al., 2016; Kittur et al., 2012; Wu et al., 2023c, 2022a). LLM chains rarely implement dynamic architectures for creative work (Yao et al., 2023; Wu et al., 2023c; Kim et al., 2023b; Zhang et al., 2023a). Crowdsourcing work applies dynamic architectures to a wider variety of creative tasks (Valentine et al., 2017; Lasecki et al., 2015; Retelny et al., 2014; Chilton et al., 2019; Kittur et al., 2011) and includes more complex dynamic workflows such as multi-level map-reduce (Kittur et al., 2011) and dynamically choosing between workflows using their maximum expected value of utility (Goto et al., 2016). Decision theoretic techniques, such as Partially Observable Markov Decision Processes, yield great efficiencies in crowdsourcing (Dai et al., 2010, 2013), but have thus far barely been considered for LLM chaining (Madaan et al., 2023a).
We recommend pursuing a broader range of creative tasks with more communicative (e.g. establishing a hierarchy of roles) and dynamic (e.g. multi-level map reduce) architectures.
8.2. Integrating the user
Although many chaining papers discuss involving humans into the chain hypothetically, few implement this user oversight (Stretcu et al., 2023; Toubal et al., 2024). Our case studies showcase one example of interactions through directly manipulable outputs, but the varied ways of incorporating the user from crowdsourcing workflows could continue to inspire similar approaches in chaining.
We recommend incorporating more approaches to user guidance from crowdsourcing, such as explorable interfaces (Willett et al., 2012; Chang et al., 2017; Cheng and Bernstein, 2015; Mohanty et al., 2019), worker-initiated revision (Salehi et al., 2017), and guided prototyping (Nebeling et al., 2016; Lasecki et al., 2015). We also recommend research into the the opportunities and limitations of user guidance, especially for creative tasks, as initial findings conflict on if using generative AI tools leads to divergent thinking (chi, 2023) or design fixation (Wadinambiarachchi et al., 2024).
8.3. Studying chains more systematically
More systematic studies on optimal chain design would be valuable. Several studies already approach questions around model choice (Madaan et al., 2023a; Du et al., 2023), incorporating roles into prompts (Wu et al., 2023a; Park et al., 2022), and using self-reflection or debate to validate outputs (Du et al., 2023; Liang et al., 2023). Although the outcomes may be different, systematic crowdsourcing studies could provide inspiration for some of these explorations. Prior crowdsourcing studies investigate the effect of context timing, architecture, and content on task output (Luther et al., 2015; Salehi et al., 2017). Others measure tradeoffs of iterative and parallel workflows (e.g., effect on latency, quality averages and variance, risk of cascading errors) in various task contexts, as well as good practices for each (e.g., appropriate number of branches or iterations and what context to share amongst branches) (Kittur et al., 2011; Chang et al., 2017; Luther et al., 2015; Huffaker et al., 2020; Little et al., 2010a; Goto et al., 2016; André et al., 2014).
By specifying a strategy of interest, the design space can orient studies for understanding optimal tactical design decisions. For example, studying debate designs could look at debate to establish adaptability for tasks with many edge cases, debate to validate outputs, or debate to encourage a diversity of creative ideas. Evaluating the use of tactic designs within the context of specific strategies will lead to more generalizable findings than evaluating solely on task performance.
We recommend systematic tests for tactical design decisions informed by prior crowdsourcing work, objectives, and strategies.
8.4. LLM workflow design processes are unstructured and design support is needed
Since LLMs are cheaper and faster to query (Terwiesch and Ulrich, [n. d.]), there are more opportunities to find optimal designs through systematic testing. While the lower barrier to entry may allow LLM chain designers to iterate more quickly, it doesn’t guarantee that they focus their attention on the right things while iterating. Since LLMs display unpredictable outputs in response to low-level changes, designers may end up fixating on minor tweaks like prompt rewordings rather than exploring the design space of strategies and tactics. To avoid potential design fixation that tunnels into technical detail rather than exploring higher-level design directions, LLM chain designers would benefit from tools that scaffold their design and iteration process. Although running an LLM chain is cheaper and faster than running a crowdsourcing workflow, the design process is unstructured. There is a need for design support.
We see many promising directions to help build chains efficiently and effectively, given that iterations are often based on extensive qualitative analysis and enlightened trial-and-error.
As there are many overlapping tactics among workflows, the flexibility and reusability of modules can support the development of future workflows (Josifoski et al., 2023). A repository of “battle-tested” workflow operators could allow sharing of quality parts for reuse in new workflows. Such a platform could include operator-level test suites to assist with validation that would monitor operator output quality in new content areas or with newly released LLMs. The platform design could draw on crowdsourcing to create such test suites. Additionally, corresponding workflow design tools could suggest different options to achieve a goal by collecting multiple approaches for sub-parts of the workflow.
Other work engineers more optimal LLM prompts by iteratively generating and scoring test prompt candidates using mathematical or natural language optimization functions (Zhou et al., 2022a; Yang et al., 2023). Future work could adapt these approaches within the context of chaining, for example to optimize for diverse responses.
Comparing many workflow designs can be difficult because of the requirement that the output of a previous step match the input of the next step. Therefore, removing or adding steps can be a non-trivial process. More formalized input and output indicators, potentially with matching prompts, could facilitate easier chain development. This approximates the approach of Flash Teams which documents the inputs and outputs of steps (Retelny et al., 2014).
Sketching aids practitioners in rapidly instantiating ideas to explore the possible solution space rather than committing too early to one design direction (Buxton, 2010; Dow et al., 2011; Yang et al., 2019). Incorporating sketching could aid LLM chain designers in arriving at better solutions, rather than getting stuck in local optima. Sketches are quick, but they also should be low-fidelity and minimal in detail. While LLMs afford quick instantiation, they produce high-fidelity output that can distract designers from their primary prototyping goals (Jiang et al., 2022; Lam et al., 2023). Opportunities to support this designing perspective include providing interfaces amenable to rapid chain construction and reconfiguration (Wu et al., 2022a; Arawjo et al., 2023) and suggesting strategies and tactics when a user may be tunneling on technical details.
We recommend continued work on chain design support tools such as repositories of workflow operators, prompting optimization methods for chaining, formalized input-output indicators, and sketching-inspired workflow design approaches.
9. Limitations
Our work addresses crowdsourcing workflows and LLM chains, but it is limited in scope beyond these fields and in discussion of broader impacts. Additionally, our recommendations are contingent upon an imperfect understanding of LLM capabilities, and we support the utility of our case studies through theory instead of a user study.
Scope. Our work targets LLM chain design by adapting crowdsourcing workflow techniques. However, there are opportunities to look beyond LLMs to multi-modal models as the workers (Gupta and Kembhavi, 2023; Surís et al., 2023; Liu et al., 2022) and beyond crowdsourcing to other task decomposition methods as a source of inspiration (Anya, 2015; Andreas et al., 2016; Mao et al., 2019).
Additionally, there may be limitations to the scope of work covered by our review of the LLM chaining literature. Since LLM chaining is a relatively new area, researchers have not converged upon established terminology as in the crowdsourcing domain. Our keyword search may not have captured some work that did not use one of our keyword search terms.
Broader impacts. We focus on the ability of workflows to serve a beneficial role in supporting various tasks. However, there are potential risks in using LLMs and crowdworkers as support tools that we do not explicitly consider. Several papers in the design space cite concerns around a loss of jobs and unintended consequences of using LLMs, especially for creative tasks (Park et al., 2022; Mirowski et al., 2023; Wu et al., 2023a). Crowdsourcing work similarly raised concerns around replacing expert workers and around labor rights for crowdworkers (Kittur et al., 2013). Broader discussions about privacy (Wu et al., 2023a; Peris et al., 2023), ownership (Kim et al., 2023b; Mirowski et al., 2023), and perceived sense of agency we also leave to future work. Finally, although workflows can improve output quality, results will still be imperfect, so errors from implemented chains could still have harmful impacts (Eggmann et al., 2023; Greshake et al., 2023).
Rapid changes of LLM capabilities. The capabilities of LLMs are rapidly changing, so researchers’ understanding of and subsequent recommendations to work with these capabilities may fall out of sync with new model developments. We recommend future work to monitor the efficacy of chain designs as new models are released, with potentially differing capabilities and failure modes.
Lack of user evaluation. We build on longstanding HCI theory and practice that demonstrates the benefit of direct manipulation interfaces (Shneiderman, 1982; Hutchins et al., 1985). However, our work does not conduct a user evaluation to explicitly test the benefits and drawbacks of our direct manipulation case study implementations. Future work may more deeply explore how users interact with LLM chains via direct manipulation and whether these interactions positively impact user control and task outcomes.
10. Conclusion
By revisiting the established field of crowdsourcing workflows, we find inspiration and direction for the nascent area of LLM chaining. To facilitate adaptation, we contribute a design space for LLM chain designers that incorporates findings from the crowdsourcing workflow and LLM chaining literatures. This design space covers the objectives supported by and the tactical building blocks of prior workflows and chains. We also surface strategies employed by designers to use tactics to support objectives. Through three case studies, we build an understanding for how crowdsourcing workflows can be adapted for LLM chains.
Finally, we synthesize our findings to discuss tactical design decisions for LLM chains through the context of their comparison with crowdsourcing workflows. We find LLM chains need additional support in subtask design and validation due in part to differences between LLM stochasticity and human diversity. On the other hand, the comparatively low cost and latency of LLM calls allows for faster iteration and for workflow techniques that were studied but never practical with human crowds. We hope these findings guide future work towards more approaches to user guidance and creative tasks, systematic discovery of good practices for tactic design, and artifacts to support the chain design process.
LLM chains present remarkable flexibility, which comes along with both promising capabilities but also substantial complexity. Our work helps LLM chain designers to navigate this vast design space by building on the established field of crowdsourcing.
Acknowledgements.
We thank Michael Bernstein and Joy Kim for providing outputs from the original Soylent and Mechanical Novel papers. We also thank the Allen Institute for Artificial Intelligence for access to Amazon Mechanical Turk resources. We thank OpenAI for offering research credits to the GPT-3.5 API. This research was supported in part by NSF award IIS-1901386, ONR grant N00014-21-1-2707, and the Allen Institute for Artificial Intelligence (AI2).References
- (1)
- chi (2023) 2023. DIS ’23: Proceedings of the 2023 ACM Designing Interactive Systems Conference (, Pittsburgh, PA, USA,). Association for Computing Machinery, New York, NY, USA.
- Agapie et al. (2015) Elena Agapie, Jaime Teevan, and Andrés Monroy-Hernández. 2015. Crowdsourcing in the field: A case study using local crowds for event reporting. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 3. 2–11.
- Ahmad et al. (2011) Salman Ahmad, Alexis Battle, Zahan Malkani, and Sepander Kamvar. 2011. The jabberwocky programming environment for structured social computing. In Proceedings of the 24th annual ACM symposium on User interface software and technology. 53–64.
- Alexander (1977) Christopher Alexander. 1977. A Pattern Language: Towns, Buildings, Construction. Oxford University Press.
- ALLEN et al. (2023) GARRETT ALLEN, GAOLE HE, and UJWAL GADIRAJU. 2023. Power-up! What Can Generative Models Do for Human Computation Workflows? (2023).
- Alshaibani et al. (2020) Abdullah Alshaibani, Sylvia Carrell, Li-Hsin Tseng, Jungmin Shin, and Alexander Quinn. 2020. Privacy-preserving face redaction using crowdsourcing. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 8. 13–22.
- Alshaibani and Quinn (2021) Abdullah Alshaibani and Alexander J Quinn. 2021. Pterodactyl: Two-Step Redaction of Images for Robust Face Deidentification. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 9. 27–34.
- Ambati et al. (2012) Vamshi Ambati, Stephan Vogel, and Jaime Carbonell. 2012. Collaborative workflow for crowdsourcing translation. In Proceedings of the ACM 2012 Conference on Computer Supported Cooperative Work. 1191–1194.
- André et al. (2014) Paul André, Robert E Kraut, and Aniket Kittur. 2014. Effects of simultaneous and sequential work structures on distributed collaborative interdependent tasks. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. 139–148.
- André et al. (2013) Paul André, Haoqi Zhang, Juho Kim, Lydia Chilton, Steven Dow, and Robert Miller. 2013. Community clustering: Leveraging an academic crowd to form coherent conference sessions. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 1. 9–16.
- Andreas et al. (2016) Jacob Andreas, Marcus Rohrbach, Trevor Darrell, and Dan Klein. 2016. Neural module networks. In Proceedings of the IEEE conference on computer vision and pattern recognition. 39–48.
- Anya (2015) Obinna Anya. 2015. Bridge the gap! What can work design in crowdwork learn from work design theories?. In Proceedings of the 18th ACM Conference on Computer Supported Cooperative Work & Social Computing. 612–627.
- Arawjo et al. (2023) Ian Arawjo, Priyan Vaithilingam, Martin Wattenberg, and Elena Glassman. 2023. ChainForge: An open-source visual programming environment for prompt engineering. In Adjunct Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. 1–3.
- Arora et al. (2022) Simran Arora, Avanika Narayan, Mayee F Chen, Laurel J Orr, Neel Guha, Kush Bhatia, Ines Chami, Frederic Sala, and Christopher Ré. 2022. Ask me anything: A simple strategy for prompting language models. arXiv preprint arXiv:2210.02441 (2022).
- Bae et al. (2022) S Sandra Bae, Clement Zheng, Mary Etta West, Ellen Yi-Luen Do, Samuel Huron, and Danielle Albers Szafir. 2022. Making data tangible: A cross-disciplinary design space for data physicalization. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems. 1–18.
- Bender et al. (2021) Emily M Bender, Timnit Gebru, Angelina McMillan-Major, and Shmargaret Shmitchell. 2021. On the dangers of stochastic parrots: Can language models be too big?. In Proceedings of the 2021 ACM conference on fairness, accountability, and transparency. 610–623.
- Bernstein et al. (2011) Michael S Bernstein, Joel Brandt, Robert C Miller, and David R Karger. 2011. Crowds in two seconds: Enabling realtime crowd-powered interfaces. In Proceedings of the 24th annual ACM symposium on User interface software and technology. 33–42.
- Bernstein et al. (2010) Michael S Bernstein, Greg Little, Robert C Miller, Björn Hartmann, Mark S Ackerman, David R Karger, David Crowell, and Katrina Panovich. 2010. Soylent: a word processor with a crowd inside. In Proceedings of the 23nd annual ACM symposium on User interface software and technology. 313–322.
- Bragg et al. (2021) Danielle Bragg, Naomi Caselli, John W Gallagher, Miriam Goldberg, Courtney J Oka, and William Thies. 2021. ASL sea battle: gamifying sign language data collection. In Proceedings of the 2021 CHI conference on human factors in computing systems. 1–13.
- Bragg et al. (2018) Jonathan Bragg, Mausam, and Daniel S. Weld. 2018. Sprout: Crowd-Powered Task Design for Crowdsourcing. Proceedings of the 31st Annual ACM Symposium on User Interface Software and Technology (2018). https://api.semanticscholar.org/CorpusID:51948377
- Bragg et al. (2013) Jonathan Bragg, Daniel Weld, et al. 2013. Crowdsourcing multi-label classification for taxonomy creation. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 1. 25–33.
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using thematic analysis in psychology. Qualitative research in psychology 3, 2 (2006), 77–101.
- Bursztyn et al. (2022) Victor S Bursztyn, David Demeter, Doug Downey, and Larry Birnbaum. 2022. Learning to Perform Complex Tasks through Compositional Fine-Tuning of Language Models. arXiv preprint arXiv:2210.12607 (2022).
- Butler et al. (2017) Crystal Butler, Stephanie Michalowicz, Lakshmi Subramanian, and Winslow Burleson. 2017. More than a Feeling: The MiFace Framework for Defining Facial Communication Mappings. In Proceedings of the 30th Annual ACM Symposium on User Interface Software and Technology. 773–786.
- Buxton (2010) Bill Buxton. 2010. Sketching User Experiences: Getting the Design Right and the Right Design. Morgan Kaufmann.
- Chang et al. (2017) Joseph Chee Chang, Saleema Amershi, and Ece Kamar. 2017. Revolt: Collaborative crowdsourcing for labeling machine learning datasets. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems. 2334–2346.
- Chen et al. (2019) Quanze Chen, Jonathan Bragg, Lydia B Chilton, and Dan S Weld. 2019. Cicero: Multi-turn, contextual argumentation for accurate crowdsourcing. In Proceedings of the 2019 chi conference on human factors in computing systems. 1–14.
- Cheng and Bernstein (2015) Justin Cheng and Michael S Bernstein. 2015. Flock: Hybrid crowd-machine learning classifiers. In Proceedings of the 18th ACM conference on computer supported cooperative work & social computing. 600–611.
- Cheng et al. (2015) Justin Cheng, Jaime Teevan, Shamsi T Iqbal, and Michael S Bernstein. 2015. Break it down: A comparison of macro-and microtasks. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems. 4061–4064.
- Chilton et al. (2014) Lydia B Chilton, Juho Kim, Paul André, Felicia Cordeiro, James A Landay, Daniel S Weld, Steven P Dow, Robert C Miller, and Haoqi Zhang. 2014. Frenzy: collaborative data organization for creating conference sessions. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. 1255–1264.
- Chilton et al. (2013) Lydia B Chilton, Greg Little, Darren Edge, Daniel S Weld, and James A Landay. 2013. Cascade: Crowdsourcing taxonomy creation. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. 1999–2008.
- Chilton et al. (2019) Lydia B Chilton, Savvas Petridis, and Maneesh Agrawala. 2019. VisiBlends: A flexible workflow for visual blends. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems. 1–14.
- Christoforaki and Ipeirotis (2014) Maria Christoforaki and Panagiotis Ipeirotis. 2014. Step: A scalable testing and evaluation platform. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 2. 41–49.
- Cobbe et al. (2021) Karl Cobbe, Vineet Kosaraju, Mohammad Bavarian, Mark Chen, Heewoo Jun, Lukasz Kaiser, Matthias Plappert, Jerry Tworek, Jacob Hilton, Reiichiro Nakano, et al. 2021. Training verifiers to solve math word problems. arXiv preprint arXiv:2110.14168 (2021).
- Creswell and Shanahan (2022) Antonia Creswell and Murray Shanahan. 2022. Faithful reasoning using large language models. arXiv preprint arXiv:2208.14271 (2022).
- Creswell et al. (2022) Antonia Creswell, Murray Shanahan, and Irina Higgins. 2022. Selection-inference: Exploiting large language models for interpretable logical reasoning. arXiv preprint arXiv:2205.09712 (2022).
- Dai et al. (2013) Peng Dai, Christopher H Lin, Daniel S Weld, et al. 2013. POMDP-based control of workflows for crowdsourcing. Artificial Intelligence 202 (2013), 52–85.
- Dai et al. (2010) Peng Dai, Daniel Weld, et al. 2010. Decision-theoretic control of crowd-sourced workflows. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 24. 1168–1174.
- Difallah et al. (2018) Djellel Difallah, Elena Filatova, and Panos Ipeirotis. 2018. Demographics and dynamics of mechanical turk workers. In Proceedings of the eleventh ACM international conference on web search and data mining. 135–143.
- Dow et al. (2011) Steven P. Dow, Alana Glassco, Jonathan Kass, Melissa Schwarz, Daniel L. Schwartz, and Scott R. Klemmer. 2011. Parallel Prototyping Leads to Better Design Results, More Divergence, and Increased Self-Efficacy. ACM Trans. Comput.-Hum. Interact. 17, 4, Article 18 (dec 2011), 24 pages. https://doi.org/10.1145/1879831.1879836
- Drapeau et al. (2016) Ryan Drapeau, Lydia Chilton, Jonathan Bragg, and Daniel Weld. 2016. Microtalk: Using argumentation to improve crowdsourcing accuracy. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 4. 32–41.
- Du et al. (2023) Yilun Du, Shuang Li, Antonio Torralba, Joshua B Tenenbaum, and Igor Mordatch. 2023. Improving Factuality and Reasoning in Language Models through Multiagent Debate. arXiv preprint arXiv:2305.14325 (2023).
- Dziri et al. (2023) Nouha Dziri, Ximing Lu, Melanie Sclar, Xiang Lorraine Li, Liwei Jian, Bill Yuchen Lin, Peter West, Chandra Bhagavatula, Ronan Le Bras, Jena D Hwang, et al. 2023. Faith and Fate: Limits of Transformers on Compositionality. arXiv preprint arXiv:2305.18654 (2023).
- Dziri et al. (2022) Nouha Dziri, Sivan Milton, Mo Yu, Osmar R Zaiane, and Siva Reddy. 2022. On the Origin of Hallucinations in Conversational Models: Is it the Datasets or the Models? ArXiv abs/2204.07931 (2022). https://api.semanticscholar.org/CorpusID:248227301
- Eggmann et al. (2023) Florin Eggmann, Roland Weiger, Nicola U Zitzmann, and Markus B Blatz. 2023. Implications of large language models such as ChatGPT for dental medicine. Journal of Esthetic and Restorative Dentistry (2023).
- Eisenstein et al. (2022) Jacob Eisenstein, Daniel Andor, Bernd Bohnet, Michael Collins, and David Mimno. 2022. Honest students from untrusted teachers: Learning an interpretable question-answering pipeline from a pretrained language model. arXiv preprint arXiv:2210.02498 (2022).
- Fleming et al. (2023) Scott L Fleming, Keith Morse, Aswathi M Kumar, Chia-Chun Chiang, Birju Patel, Emma P Brunskill, and Nigam Shah. 2023. Assessing the Potential of USMLE-Like Exam Questions Generated by GPT-4. medRxiv (2023), 2023–04.
- Fricke (2018) Suzanne Fricke. 2018. Semantic scholar. Journal of the Medical Library Association: JMLA 106, 1 (2018), 145.
- Gamma et al. (1994) Erich Gamma, Richard Helm, Ralph Johnson, and John Vlissides. 1994. Design Patterns Elements of Reusable Object-Oriented Software. Addison-Wesley.
- Gero et al. (2023) Zelalem Gero, Chandan Singh, Hao Cheng, Tristan Naumann, Michel Galley, Jianfeng Gao, and Hoifung Poon. 2023. Self-Verification Improves Few-Shot Clinical Information Extraction. arXiv preprint arXiv:2306.00024 (2023).
- Gordon et al. (2022) Mitchell L. Gordon, Michelle S. Lam, Joon Sung Park, Kayur Patel, Jeff Hancock, Tatsunori Hashimoto, and Michael S. Bernstein. 2022. Jury Learning: Integrating Dissenting Voices into Machine Learning Models. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (, New Orleans, LA, USA,) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 115, 19 pages. https://doi.org/10.1145/3491102.3502004
- Goto et al. (2016) Shinsuke Goto, Toru Ishida, and Donghui Lin. 2016. Understanding crowdsourcing workflow: modeling and optimizing iterative and parallel processes. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 4. 52–58.
- Gouravajhala et al. (2018) Sai Gouravajhala, Jinyeong Yim, Karthik Desingh, Yanda Huang, Odest Jenkins, and Walter Lasecki. 2018. Eureca: Enhanced understanding of real environments via crowd assistance. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 6. 31–40.
- Greshake et al. (2023) Kai Greshake, Sahar Abdelnabi, Shailesh Mishra, Christoph Endres, Thorsten Holz, and Mario Fritz. 2023. Not what you’ve signed up for: Compromising Real-World LLM-Integrated Applications with Indirect Prompt Injection. arXiv preprint arXiv:2302.12173 (2023).
- Guo et al. (2016) Anhong Guo, Xiang’Anthony’ Chen, Haoran Qi, Samuel White, Suman Ghosh, Chieko Asakawa, and Jeffrey P Bigham. 2016. Vizlens: A robust and interactive screen reader for interfaces in the real world. In Proceedings of the 29th annual symposium on user interface software and technology. 651–664.
- Guo et al. (2017) Anhong Guo, Jeeeun Kim, Xiang’Anthony’ Chen, Tom Yeh, Scott E Hudson, Jennifer Mankoff, and Jeffrey P Bigham. 2017. Facade: Auto-generating tactile interfaces to appliances. In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems. 5826–5838.
- Guo et al. (2019) Anhong Guo, Junhan Kong, Michael Rivera, Frank F Xu, and Jeffrey P Bigham. 2019. Statelens: A reverse engineering solution for making existing dynamic touchscreens accessible. In Proceedings of the 32nd Annual ACM Symposium on User Interface Software and Technology. 371–385.
- Gupta and Kembhavi (2023) Tanmay Gupta and Aniruddha Kembhavi. 2023. Visual programming: Compositional visual reasoning without training. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 14953–14962.
- Hahn et al. (2016) Nathan Hahn, Joseph Chang, Ji Eun Kim, and Aniket Kittur. 2016. The Knowledge Accelerator: Big picture thinking in small pieces. In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems. 2258–2270.
- Hata et al. (2017) Kenji Hata, Ranjay Krishna, Li Fei-Fei, and Michael S Bernstein. 2017. A glimpse far into the future: Understanding long-term crowd worker quality. In Proceedings of the 2017 ACM Conference on Computer Supported Cooperative Work and Social Computing. 889–901.
- Holtzman et al. (2021) Ari Holtzman, Peter West, Vered Shwartz, Yejin Choi, and Luke Zettlemoyer. 2021. Surface form competition: Why the highest probability answer isn’t always right. arXiv preprint arXiv:2104.08315 (2021).
- Horvitz (1999) Eric Horvitz. 1999. Principles of mixed-initiative user interfaces. In Proceedings of the SIGCHI conference on Human Factors in Computing Systems. 159–166.
- Hsieh et al. (2023) Cheng-Yu Hsieh, Si-An Chen, Chun-Liang Li, Yasuhisa Fujii, Alexander Ratner, Chen-Yu Lee, Ranjay Krishna, and Tomas Pfister. 2023. Tool documentation enables zero-shot tool-usage with large language models. arXiv preprint arXiv:2308.00675 (2023).
- Huang et al. (2023a) Qing Huang, Yishun Wu, Zhenchang Xing, He Jiang, Yu Cheng, and Huan Jin. 2023a. Adaptive Intellect Unleashed: The Feasibility of Knowledge Transfer in Large Language Models. arXiv preprint arXiv:2308.04788 (2023).
- Huang et al. (2023b) Qing Huang, Jiahui Zhu, Zhilong Li, Zhenchang Xing, Changjing Wang, and Xiwei Xu. 2023b. Pcr-chain: Partial code reuse assisted by hierarchical chaining of prompts on frozen copilot. In 2023 IEEE/ACM 45th International Conference on Software Engineering: Companion Proceedings (ICSE-Companion). IEEE, 1–5.
- Huang et al. (2023c) Qing Huang, Zhou Zou, Zhenchang Xing, Zhenkang Zuo, Xiwei Xu, and Qinghua Lu. 2023c. AI Chain on Large Language Model for Unsupervised Control Flow Graph Generation for Statically-Typed Partial Code. arXiv preprint arXiv:2306.00757 (2023).
- Huang et al. (2018) Ting-Hao Huang, Joseph Chee Chang, and Jeffrey P Bigham. 2018. Evorus: A crowd-powered conversational assistant built to automate itself over time. In Proceedings of the 2018 CHI conference on human factors in computing systems. 1–13.
- Huang et al. (2015) Ting-Hao Huang, Walter Lasecki, and Jeffrey Bigham. 2015. Guardian: A crowd-powered spoken dialog system for web apis. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 3. 62–71.
- Huang et al. (2017) Yi-Ching Huang, Jiunn-Chia Huang, Hao-Chuan Wang, and Jane Hsu. 2017. Supporting ESL writing by prompting crowdsourced structural feedback. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 5. 71–78.
- Huffaker et al. (2020) Jordan S Huffaker, Jonathan K Kummerfeld, Walter S Lasecki, and Mark S Ackerman. 2020. Crowdsourced detection of emotionally manipulative language. In Proceedings of the 2020 CHI conference on human factors in computing systems. 1–14.
- Hutchins et al. (1985) Edwin L Hutchins, James D Hollan, and Donald A Norman. 1985. Direct manipulation interfaces. Human–computer interaction 1, 4 (1985), 311–338.
- Jiang et al. (2022) Ellen Jiang, Kristen Olson, Edwin Toh, Alejandra Molina, Aaron Donsbach, Michael Terry, and Carrie J Cai. 2022. PromptMaker: Prompt-Based Prototyping with Large Language Models. In Extended Abstracts of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI EA ’22). Association for Computing Machinery, New York, NY, USA, Article 35, 8 pages. https://doi.org/10.1145/3491101.3503564
- Josifoski et al. (2023) Martin Josifoski, Lars Klein, Maxime Peyrard, Yifei Li, Saibo Geng, Julian Paul Schnitzler, Yuxing Yao, Jiheng Wei, Debjit Paul, and Robert West. 2023. Flows: Building blocks of reasoning and collaborating ai. arXiv preprint arXiv:2308.01285 (2023).
- Kaur et al. (2017) Harmanpreet Kaur, Mitchell Gordon, Yiwei Yang, Jeffrey Bigham, Jaime Teevan, Ece Kamar, and Walter Lasecki. 2017. Crowdmask: Using crowds to preserve privacy in crowd-powered systems via progressive filtering. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 5. 89–98.
- Kim and Monroy-Hernandez (2016) Joy Kim and Andres Monroy-Hernandez. 2016. Storia: Summarizing social media content based on narrative theory using crowdsourcing. In Proceedings of the 19th ACM Conference on Computer-Supported Cooperative Work & Social Computing. 1018–1027.
- Kim et al. (2014) Juho Kim, Phu Tran Nguyen, Sarah Weir, Philip J Guo, Robert C Miller, and Krzysztof Z Gajos. 2014. Crowdsourcing step-by-step information extraction to enhance existing how-to videos. In Proceedings of the SIGCHI conference on human factors in computing systems. 4017–4026.
- Kim et al. (2017) Joy Kim, Sarah Sterman, Allegra Argent Beal Cohen, and Michael S Bernstein. 2017. Mechanical novel: Crowdsourcing complex work through reflection and revision. In Proceedings of the 2017 acm conference on computer supported cooperative work and social computing. 233–245.
- Kim et al. (2023b) Jeongyeon Kim, Sangho Suh, Lydia B Chilton, and Haijun Xia. 2023b. Metaphorian: Leveraging Large Language Models to Support Extended Metaphor Creation for Science Writing. In Proceedings of the 2023 ACM Designing Interactive Systems Conference. 115–135.
- Kim et al. (2023a) Tae Soo Kim, Yoonjoo Lee, Minsuk Chang, and Juho Kim. 2023a. Cells, generators, and lenses: Design framework for object-oriented interaction with large language models. In Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology. 1–18.
- Kittur et al. (2012) Aniket Kittur, Susheel Khamkar, Paul André, and Robert Kraut. 2012. CrowdWeaver: visually managing complex crowd work. In Proceedings of the ACM 2012 Conference on Computer Supported Cooperative Work. 1033–1036.
- Kittur et al. (2013) Aniket Kittur, Jeffrey V Nickerson, Michael Bernstein, Elizabeth Gerber, Aaron Shaw, John Zimmerman, Matt Lease, and John Horton. 2013. The future of crowd work. In Proceedings of the 2013 conference on Computer supported cooperative work. 1301–1318.
- Kittur et al. (2011) Aniket Kittur, Boris Smus, Susheel Khamkar, and Robert E Kraut. 2011. Crowdforge: Crowdsourcing complex work. In Proceedings of the 24th annual ACM symposium on User interface software and technology. 43–52.
- Kobayashi et al. (2018) Masaki Kobayashi, Hiromi Morita, Masaki Matsubara, Nobuyuki Shimizu, and Atsuyuki Morishima. 2018. An empirical study on short-and long-term effects of self-correction in crowdsourced microtasks. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 6. 79–87.
- Krishna (2019) Ranjay Krishna. 2019. EasyTurk: A Wrapper for Custom AMT Tasks. https://github.com/ranjaykrishna/easyturk.
- Krizhevsky et al. (2009) Alex Krizhevsky, Geoffrey Hinton, et al. 2009. Learning multiple layers of features from tiny images. (2009).
- Kulkarni et al. (2012) Anand Kulkarni, Matthew Can, and Björn Hartmann. 2012. Collaboratively crowdsourcing workflows with turkomatic. In Proceedings of the acm 2012 conference on computer supported cooperative work. 1003–1012.
- Kulkarni et al. (2014) Anand Kulkarni, Prayag Narula, David Rolnitzky, and Nathan Kontny. 2014. Wish: Amplifying creative ability with expert crowds. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 2. 112–120.
- Lai et al. (2023) Vivian Lai, Chacha Chen, Alison Smith-Renner, Q Vera Liao, and Chenhao Tan. 2023. Towards a Science of Human-AI Decision Making: An Overview of Design Space in Empirical Human-Subject Studies. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency. 1369–1385.
- Lam et al. (2023) Michelle S. Lam, Zixian Ma, Anne Li, Izequiel Freitas, Dakuo Wang, James A. Landay, and Michael S. Bernstein. 2023. Model Sketching: Centering Concepts in Early-Stage Machine Learning Model Design. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 741, 24 pages. https://doi.org/10.1145/3544548.3581290
- LangChain (2022) LangChain. 2022. LangChain. (2022). https://www.langchain.com/
- Lasecki et al. (2015) Walter S Lasecki, Juho Kim, Nick Rafter, Onkur Sen, Jeffrey P Bigham, and Michael S Bernstein. 2015. Apparition: Crowdsourced user interfaces that come to life as you sketch them. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems. 1925–1934.
- Lasecki et al. (2013) Walter S Lasecki, Rachel Wesley, Jeffrey Nichols, Anand Kulkarni, James F Allen, and Jeffrey P Bigham. 2013. Chorus: a crowd-powered conversational assistant. In Proceedings of the 26th annual ACM symposium on User interface software and technology. 151–162.
- LaToza et al. (2014) Thomas D LaToza, W Ben Towne, Christian M Adriano, and André Van Der Hoek. 2014. Microtask programming: Building software with a crowd. In Proceedings of the 27th annual ACM symposium on User interface software and technology. 43–54.
- Lee et al. (2017) Sang Won Lee, Yujin Zhang, Isabelle Wong, Yiwei Yang, Stephanie D O’Keefe, and Walter S Lasecki. 2017. Sketchexpress: Remixing animations for more effective crowd-powered prototyping of interactive interfaces. In Proceedings of the 30th Annual ACM Symposium on User Interface Software and Technology. 817–828.
- Levine et al. (2022) Yoav Levine, Itay Dalmedigos, Ori Ram, Yoel Zeldes, Daniel Jannai, Dor Muhlgay, Yoni Osin, Opher Lieber, Barak Lenz, Shai Shalev-Shwartz, et al. 2022. Standing on the shoulders of giant frozen language models. arXiv preprint arXiv:2204.10019 (2022).
- Li et al. (2023) Cheng Li, Mingyang Zhang, Qiaozhu Mei, Yaqing Wang, Spurthi Amba Hombaiah, Yi Liang, and Michael Bendersky. 2023. Teach LLMs to Personalize–An Approach inspired by Writing Education. arXiv preprint arXiv:2308.07968 (2023).
- Li (2024) Jiyi Li. 2024. A Comparative Study on Annotation Quality of Crowdsourcing and LLM via Label Aggregation. CoRR abs/2401.09760 (2024). https://doi.org/10.48550/ARXIV.2401.09760 arXiv:2401.09760
- Liang et al. (2023) Tian Liang, Zhiwei He, Wenxiang Jiao, Xing Wang, Yan Wang, Rui Wang, Yujiu Yang, Zhaopeng Tu, and Shuming Shi. 2023. Encouraging Divergent Thinking in Large Language Models through Multi-Agent Debate. arXiv preprint arXiv:2305.19118 (2023).
- Liem et al. (2011) Beatrice Liem, Haoqi Zhang, and Yiling Chen. 2011. An iterative dual pathway structure for speech-to-text transcription. (2011).
- Lin et al. (2014a) Chi-Chin Lin, Yi-Ching Huang, and Jane Yung-jen Hsu. 2014a. Crowdsourced explanations for humorous internet memes based on linguistic theories. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 2. 143–150.
- Lin et al. (2014b) Tsung-Yi Lin, Michael Maire, Serge Belongie, James Hays, Pietro Perona, Deva Ramanan, Piotr Dollár, and C Lawrence Zitnick. 2014b. Microsoft coco: Common objects in context. In Computer Vision–ECCV 2014: 13th European Conference, Zurich, Switzerland, September 6-12, 2014, Proceedings, Part V 13. Springer, 740–755.
- Little et al. (2010a) Greg Little, Lydia B Chilton, Max Goldman, and Robert C Miller. 2010a. Exploring iterative and parallel human computation processes. In Proceedings of the ACM SIGKDD workshop on human computation. 68–76.
- Little et al. (2010b) Greg Little, Lydia B Chilton, Max Goldman, and Robert C Miller. 2010b. Turkit: human computation algorithms on mechanical turk. In Proceedings of the 23nd annual ACM symposium on User interface software and technology. 57–66.
- Liu et al. (2016) Angli Liu, Stephen Soderland, Jonathan Bragg, C. H. Lin, Xiao Ling, and Daniel S. Weld. 2016. Effective Crowd Annotation for Relation Extraction. In North American Chapter of the Association for Computational Linguistics. https://api.semanticscholar.org/CorpusID:10705630
- Liu et al. (2018) Ching Liu, Juho Kim, and Hao-Chuan Wang. 2018. ConceptScape: Collaborative concept mapping for video learning. In Proceedings of the 2018 CHI conference on human factors in computing systems. 1–12.
- Liu et al. (2023) Ryan Liu, Howard Yen, Raja Marjieh, Thomas L Griffiths, and Ranjay Krishna. 2023. Improving interpersonal communication by simulating audiences with language models. arXiv preprint arXiv:2311.00687 (2023).
- Liu et al. (2022) Vivian Liu, Han Qiao, and Lydia Chilton. 2022. Opal: Multimodal image generation for news illustration. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology. 1–17.
- Long et al. (2023) Tao Long, Dorothy Zhang, Grace Li, Batool Taraif, Samia Menon, Kynnedy Simone Smith, Sitong Wang, Katy Ilonka Gero, and Lydia B Chilton. 2023. Tweetorial Hooks: Generative AI Tools to Motivate Science on Social Media. arXiv preprint arXiv:2305.12265 (2023).
- Luther et al. (2015) Kurt Luther, Nathan Hahn, Steven Dow, and Aniket Kittur. 2015. Crowdlines: Supporting synthesis of diverse information sources through crowdsourced outlines. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 3. 110–119.
- MacNeil et al. (2023) Stephen MacNeil, Andrew Tran, Joanne Kim, Ziheng Huang, Seth Bernstein, and Dan Mogil. 2023. Prompt Middleware: Mapping Prompts for Large Language Models to UI Affordances. arXiv preprint arXiv:2307.01142 (2023).
- Madaan et al. (2023a) Aman Madaan, Pranjal Aggarwal, Ankit Anand, Srividya Pranavi Potharaju, Swaroop Mishra, Pei Zhou, Aditya Gupta, Dheeraj Rajagopal, Karthik Kappaganthu, Yiming Yang, Shyam Upadhyay, Mausam, and Manaal Faruqui. 2023a. AutoMix: Automatically Mixing Language Models. https://api.semanticscholar.org/CorpusID:264306222
- Madaan et al. (2023b) Aman Madaan, Niket Tandon, Prakhar Gupta, Skyler Hallinan, Luyu Gao, Sarah Wiegreffe, Uri Alon, Nouha Dziri, Shrimai Prabhumoye, Yiming Yang, et al. 2023b. Self-refine: Iterative refinement with self-feedback. arXiv preprint arXiv:2303.17651 (2023).
- Mahyar et al. (2018) Narges Mahyar, Michael R James, Michelle M Ng, Reginald A Wu, and Steven P Dow. 2018. CommunityCrit: inviting the public to improve and evaluate urban design ideas through micro-activities. In Proceedings of the 2018 CHI Conference on Human Factors in Computing Systems. 1–14.
- Mao et al. (2019) Jiayuan Mao, Chuang Gan, Pushmeet Kohli, Joshua B Tenenbaum, and Jiajun Wu. 2019. The neuro-symbolic concept learner: Interpreting scenes, words, and sentences from natural supervision. arXiv preprint arXiv:1904.12584 (2019).
- Marcus et al. (2011a) Adam Marcus, Eugene Wu, David Karger, Samuel Madden, and Robert Miller. 2011a. Human-powered sorts and joins. The 38th International Conference on Very Large Data Bases (2011).
- Marcus et al. (2011b) Adam Marcus, Eugene Wu, David R Karger, Samuel Madden, and Robert C Miller. 2011b. Demonstration of qurk: a query processor for humanoperators. In Proceedings of the 2011 ACM SIGMOD International Conference on Management of data. 1315–1318.
- Maynez et al. (2020) Joshua Maynez, Shashi Narayan, Bernd Bohnet, and Ryan McDonald. 2020. On faithfulness and factuality in abstractive summarization. arXiv preprint arXiv:2005.00661 (2020).
- McDonnell et al. (2016) Tyler McDonnell, Matthew Lease, Mucahid Kutlu, and Tamer Elsayed. 2016. Why is that relevant? collecting annotator rationales for relevance judgments. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 4. 139–148.
- Mirowski et al. (2023) Piotr Mirowski, Kory W Mathewson, Jaylen Pittman, and Richard Evans. 2023. Co-Writing Screenplays and Theatre Scripts with Language Models: Evaluation by Industry Professionals. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–34.
- Mohanty et al. (2019) Vikram Mohanty, Kareem Abdol-Hamid, Courtney Ebersohl, and Kurt Luther. 2019. Second opinion: Supporting last-mile person identification with crowdsourcing and face recognition. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 7. 86–96.
- Nair et al. (2023a) Varun Nair, Elliot Schumacher, and Anitha Kannan. 2023a. Generating medically-accurate summaries of patient-provider dialogue: A multi-stage approach using large language models. arXiv preprint arXiv:2305.05982 (2023).
- Nair et al. (2023b) Varun Nair, Elliot Schumacher, Geoffrey Tso, and Anitha Kannan. 2023b. DERA: enhancing large language model completions with dialog-enabled resolving agents. arXiv preprint arXiv:2303.17071 (2023).
- Narayan et al. (2018) Shashi Narayan, Shay B Cohen, and Mirella Lapata. 2018. Don’t give me the details, just the summary! topic-aware convolutional neural networks for extreme summarization. arXiv preprint arXiv:1808.08745 (2018).
- Nebeling et al. (2016) Michael Nebeling, Alexandra To, Anhong Guo, Adrian A de Freitas, Jaime Teevan, Steven P Dow, and Jeffrey P Bigham. 2016. WearWrite: Crowd-assisted writing from smartwatches. In Proceedings of the 2016 CHI conference on human factors in computing systems. 3834–3846.
- Nie et al. (2022) Feng Nie, Meixi Chen, Zhirui Zhang, and Xu Cheng. 2022. Improving few-shot performance of language models via nearest neighbor calibration. arXiv preprint arXiv:2212.02216 (2022).
- Ouyang et al. (2022) Long Ouyang, Jeffrey Wu, Xu Jiang, Diogo Almeida, Carroll Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, et al. 2022. Training language models to follow instructions with human feedback. Advances in Neural Information Processing Systems 35 (2022), 27730–27744.
- Parameswaran et al. (2023) Aditya G Parameswaran, Shreya Shankar, Parth Asawa, Naman Jain, and Yujie Wang. 2023. Revisiting Prompt Engineering via Declarative Crowdsourcing. arXiv preprint arXiv:2308.03854 (2023).
- Park et al. (2022) Joon Sung Park, Lindsay Popowski, Carrie Cai, Meredith Ringel Morris, Percy Liang, and Michael S Bernstein. 2022. Social simulacra: Creating populated prototypes for social computing systems. In Proceedings of the 35th Annual ACM Symposium on User Interface Software and Technology. 1–18.
- Pavel et al. (2014) Amy Pavel, Colorado Reed, Björn Hartmann, and Maneesh Agrawala. 2014. Video digests: a browsable, skimmable format for informational lecture videos.. In UIST, Vol. 10. Citeseer, 2642918–2647400.
- Peris et al. (2023) Charith Peris, Christophe Dupuy, Jimit Majmudar, Rahil Parikh, Sami Smaili, Richard Zemel, and Rahul Gupta. 2023. Privacy in the Time of Language Models. In Proceedings of the Sixteenth ACM International Conference on Web Search and Data Mining. 1291–1292.
- Pietrowicz et al. (2013) Mary Pietrowicz, Danish Chopra, Amin Sadeghi, Puneet Chandra, Brian Bailey, and Karrie Karahalios. 2013. CrowdBand: An Automated Crowdsourcing Sound Composition System. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 1. 121–129.
- Quattoni and Torralba (2009) Ariadna Quattoni and Antonio Torralba. 2009. Recognizing indoor scenes. In 2009 IEEE conference on computer vision and pattern recognition. IEEE, 413–420.
- Reppert et al. (2023) Justin Reppert, Ben Rachbach, Charlie George, Luke Stebbing Jungwon Byun, Maggie Appleton, and Andreas Stuhlmüller. 2023. Iterated Decomposition: Improving Science Q&A by Supervising Reasoning Processes. arXiv preprint arXiv:2301.01751 (2023).
- Retelny et al. (2014) Daniela Retelny, Sébastien Robaszkiewicz, Alexandra To, Walter S Lasecki, Jay Patel, Negar Rahmati, Tulsee Doshi, Melissa Valentine, and Michael S Bernstein. 2014. Expert crowdsourcing with flash teams. In Proceedings of the 27th annual ACM symposium on User interface software and technology. 75–85.
- Salehi et al. (2017) Niloufar Salehi, Jaime Teevan, Shamsi Iqbal, and Ece Kamar. 2017. Communicating context to the crowd for complex writing tasks. In Proceedings of the 2017 ACM conference on computer supported cooperative work and social computing. 1890–1901.
- Salisbury et al. (2017) Elliot Salisbury, Ece Kamar, and Meredith Morris. 2017. Toward scalable social alt text: Conversational crowdsourcing as a tool for refining vision-to-language technology for the blind. In Proceedings of the AAAI Conference on Human Computation and Crowdsourcing, Vol. 5. 147–156.
- Schick et al. (2022) Timo Schick, Jane Dwivedi-Yu, Zhengbao Jiang, Fabio Petroni, Patrick Lewis, Gautier Izacard, Qingfei You, Christoforos Nalmpantis, Edouard Grave, and Sebastian Riedel. 2022. Peer: A collaborative language model. arXiv preprint arXiv:2208.11663 (2022).
- Shi et al. (2021) Yang Shi, Xingyu Lan, Jingwen Li, Zhaorui Li, and Nan Cao. 2021. Communicating with motion: A design space for animated visual narratives in data videos. In Proceedings of the 2021 CHI conference on human factors in computing systems. 1–13.
- Shneiderman (1982) Ben Shneiderman. 1982. The future of interactive systems and the emergence of direct manipulation. Behaviour & Information Technology 1, 3 (1982), 237–256.
- Sorensen et al. (2024) Taylor Sorensen, Jared Moore, Jillian Fisher, Mitchell Gordon, Niloofar Mireshghallah, Christopher Michael Rytting, Andre Ye, Liwei Jiang, Ximing Lu, Nouha Dziri, et al. 2024. A Roadmap to Pluralistic Alignment. arXiv preprint arXiv:2402.05070 (2024).
- Stretcu et al. (2023) Otilia Stretcu, Edward Vendrow, Kenji Hata, Krishnamurthy Viswanathan, Vittorio Ferrari, Sasan Tavakkol, Wenlei Zhou, Aditya Avinash, Emming Luo, Neil Gordon Alldrin, MohammadHossein Bateni, Gabriel Berger, Andrew Bunner, Chun-Ta Lu, Javier Rey, Giulia DeSalvo, Ranjay Krishna, and Ariel Fuxman. 2023. Agile Modeling: From Concept to Classifier in Minutes. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 22323–22334.
- Surís et al. (2023) Dídac Surís, Sachit Menon, and Carl Vondrick. 2023. Vipergpt: Visual inference via python execution for reasoning. arXiv preprint arXiv:2303.08128 (2023).
- Teevan et al. (2016) Jaime Teevan, Shamsi T Iqbal, and Curtis Von Veh. 2016. Supporting collaborative writing with microtasks. In Proceedings of the 2016 CHI conference on human factors in computing systems. 2657–2668.
- Terwiesch and Ulrich ([n. d.]) Christian Terwiesch and Karl Ulrich. [n. d.]. M.B.A. Students vs. ChatGPT: Who Comes Up With More Innovative Ideas? The Wall Street Journal ([n. d.]). https://www.wsj.com/tech/ai/mba-students-vs-chatgpt-innovation-679edf3b
- Toubal et al. (2024) Imad Eddine Toubal, Aditya Avinash, Neil Gordon Alldrin, Jan Dlabal, Wenlei Zhou, Enming Luo, Otilia Stretcu, Hao Xiong, Chun-Ta Lu, Howard Zhou, Ranjay Krishna, Ariel Fuxman, and Tom Duerig. 2024. Modeling Collaborator: Enabling Subjective Vision Classification With Minimal Human Effort via LLM Tool-Use. Proceedings of the IEEE/CVF International Conference on Computer Vision (2024).
- Valentine et al. (2017) Melissa A Valentine, Daniela Retelny, Alexandra To, Negar Rahmati, Tulsee Doshi, and Michael S Bernstein. 2017. Flash organizations: Crowdsourcing complex work by structuring crowds as organizations. In Proceedings of the 2017 CHI conference on human factors in computing systems. 3523–3537.
- Wadinambiarachchi et al. (2024) Samangi Wadinambiarachchi, Ryan M Kelly, Saumya Pareek, Qiushi Zhou, and Eduardo Velloso. 2024. The Effects of Generative AI on Design Fixation and Divergent Thinking. arXiv preprint arXiv:2403.11164 (2024).
- Wang et al. (2024) Angelina Wang, Jamie Morgenstern, and John P Dickerson. 2024. Large language models cannot replace human participants because they cannot portray identity groups. arXiv preprint arXiv:2402.01908 (2024).
- Wang et al. (2022) Boshi Wang, Xiang Deng, and Huan Sun. 2022. Iteratively prompt pre-trained language models for chain of thought. arXiv preprint arXiv:2203.08383 (2022).
- Wang et al. (2023) Sitong Wang, Savvas Petridis, Taeahn Kwon, Xiaojuan Ma, and Lydia B Chilton. 2023. PopBlends: Strategies for conceptual blending with large language models. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–19.
- Wei et al. (2022) Jason Wei, Xuezhi Wang, Dale Schuurmans, Maarten Bosma, Fei Xia, Ed Chi, Quoc V Le, Denny Zhou, et al. 2022. Chain-of-thought prompting elicits reasoning in large language models. Advances in Neural Information Processing Systems 35 (2022), 24824–24837.
- Weir et al. (2015) Sarah Weir, Juho Kim, Krzysztof Z Gajos, and Robert C Miller. 2015. Learnersourcing Subgoal Labels for How-to Videos. 405–416. DOI: https://doi. org/10.1145/2675133.2675219 (2015).
- Willett et al. (2012) Wesley Willett, Jeffrey Heer, and Maneesh Agrawala. 2012. Strategies for crowdsourcing social data analysis. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. 227–236.
- Wu et al. (2023a) Qingyun Wu, Gagan Bansal, Jieyu Zhang, Yiran Wu, Shaokun Zhang, Erkang Zhu, Beibin Li, Li Jiang, Xiaoyun Zhang, and Chi Wang. 2023a. AutoGen: Enabling Next-Gen LLM Applications via Multi-Agent Conversation Framework. arXiv preprint arXiv:2308.08155 (2023).
- Wu et al. (2022a) Tongshuang Wu, Ellen Jiang, Aaron Donsbach, Jeff Gray, Alejandra Molina, Michael Terry, and Carrie J Cai. 2022a. Promptchainer: Chaining large language model prompts through visual programming. In CHI Conference on Human Factors in Computing Systems Extended Abstracts. 1–10.
- Wu et al. (2022b) Tongshuang Wu, Michael Terry, and Carrie Jun Cai. 2022b. Ai chains: Transparent and controllable human-ai interaction by chaining large language model prompts. In Proceedings of the 2022 CHI conference on human factors in computing systems. 1–22.
- Wu et al. (2023c) Tongshuang Wu, Haiyi Zhu, Maya Albayrak, Alexis Axon, Amanda Bertsch, Wenxing Deng, Ziqi Ding, Bill Guo, Sireesh Gururaja, Tzu-Sheng Kuo, et al. 2023c. LLMs as Workers in Human-Computational Algorithms? Replicating Crowdsourcing Pipelines with LLMs. arXiv preprint arXiv:2307.10168 (2023).
- Wu et al. (2021) Tongshuang Sherry Wu, Michael Terry, and Carrie J. Cai. 2021. AI Chains: Transparent and Controllable Human-AI Interaction by Chaining Large Language Model Prompts. Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (2021). https://api.semanticscholar.org/CorpusID:238353829
- Wu et al. (2023b) Yiran Wu, Feiran Jia, Shaokun Zhang, Qingyun Wu, Hangyu Li, Erkang Zhu, Yue Wang, Yin Tat Lee, Richard Peng, and Chi Wang. 2023b. An Empirical Study on Challenging Math Problem Solving with GPT-4. arXiv preprint arXiv:2306.01337 (2023).
- Xie et al. (2023) Yuxi Xie, Kenji Kawaguchi, Yiran Zhao, Xu Zhao, Min-Yen Kan, Junxian He, and Qizhe Xie. 2023. Decomposition enhances reasoning via self-evaluation guided decoding. arXiv preprint arXiv:2305.00633 (2023).
- Yang et al. (2023) Chengrun Yang, Xuezhi Wang, Yifeng Lu, Hanxiao Liu, Quoc V Le, Denny Zhou, and Xinyun Chen. 2023. Large language models as optimizers. arXiv preprint arXiv:2309.03409 (2023).
- Yang et al. (2022) Kevin Yang, Nanyun Peng, Yuandong Tian, and Dan Klein. 2022. Re3: Generating longer stories with recursive reprompting and revision. arXiv preprint arXiv:2210.06774 (2022).
- Yang et al. (2019) Qian Yang, Justin Cranshaw, Saleema Amershi, Shamsi T. Iqbal, and Jaime Teevan. 2019. Sketching NLP: A Case Study of Exploring the Right Things To Design with Language Intelligence. In Proceedings of the 2019 CHI Conference on Human Factors in Computing Systems (Glasgow, Scotland Uk) (CHI ’19). Association for Computing Machinery, New York, NY, USA, 1–12. https://doi.org/10.1145/3290605.3300415
- Yao et al. (2023) Shunyu Yao, Dian Yu, Jeffrey Zhao, Izhak Shafran, Thomas L Griffiths, Yuan Cao, and Karthik Narasimhan. 2023. Tree of thoughts: Deliberate problem solving with large language models. arXiv preprint arXiv:2305.10601 (2023).
- Zaidan and Callison-Burch (2011) Omar Zaidan and Chris Callison-Burch. 2011. Crowdsourcing translation: Professional quality from non-professionals. In Proceedings of the 49th annual meeting of the association for computational linguistics: human language technologies. 1220–1229.
- Zamfirescu-Pereira et al. (2023) JD Zamfirescu-Pereira, Richmond Y Wong, Bjoern Hartmann, and Qian Yang. 2023. Why Johnny can’t prompt: how non-AI experts try (and fail) to design LLM prompts. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. 1–21.
- Zhang et al. (2012) Haoqi Zhang, Edith Law, Rob Miller, Krzysztof Gajos, David Parkes, and Eric Horvitz. 2012. Human computation tasks with global constraints. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems. 217–226.
- Zhang et al. (2023b) Jieyu Zhang, Ranjay Krishna, Ahmed H Awadallah, and Chi Wang. 2023b. EcoAssistant: Using LLM Assistant More Affordably and Accurately. arXiv preprint arXiv:2310.03046 (2023).
- Zhang et al. (2024) Shaokun Zhang, Jieyu Zhang, Jiale Liu, Linxin Song, Chi Wang, Ranjay Krishna, and Qingyun Wu. 2024. Training Language Model Agents without Modifying Language Models. International Conference on Machine Learning (2024).
- Zhang et al. (2023a) Zheng Zhang, Jie Gao, Ranjodh Singh Dhaliwal, and Toby Jia-Jun Li. 2023a. VISAR: A Human-AI Argumentative Writing Assistant with Visual Programming and Rapid Draft Prototyping. arXiv preprint arXiv:2304.07810 (2023).
- Zhang et al. (2022) Zhuosheng Zhang, Aston Zhang, Mu Li, and Alex Smola. 2022. Automatic chain of thought prompting in large language models. arXiv preprint arXiv:2210.03493 (2022).
- Zharikova et al. (2023) Diliara Zharikova, Daniel Kornev, Fedor Ignatov, Maxim Talimanchuk, Dmitry Evseev, Ksenya Petukhova, Veronika Smilga, Dmitry Karpov, Yana Shishkina, Dmitry Kosenko, et al. 2023. DeepPavlov Dream: Platform for Building Generative AI Assistants. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 3: System Demonstrations). 599–607.
- Zhou et al. (2022b) Denny Zhou, Nathanael Schärli, Le Hou, Jason Wei, Nathan Scales, Xuezhi Wang, Dale Schuurmans, Claire Cui, Olivier Bousquet, Quoc Le, et al. 2022b. Least-to-most prompting enables complex reasoning in large language models. arXiv preprint arXiv:2205.10625 (2022).
- Zhou et al. (2022a) Yongchao Zhou, Andrei Ioan Muresanu, Ziwen Han, Keiran Paster, Silviu Pitis, Harris Chan, and Jimmy Ba. 2022a. Large language models are human-level prompt engineers. arXiv preprint arXiv:2211.01910 (2022).